### v2

# One Model to Rule them All: Towards Universal Segmentation for Medical Images with Text Prompts

Abstract
Abstract. In this study, we focus on building up a model that aims to Segment Anything in medical scenarios, driven by Text prompts, termed as SAT. Our main contributions are three folds: (i) for dataset construction, we combine multiple knowledge sources to construct the first multi-modal knowledge tree on human anatomy, including 6502 anatomical terminologies; Then we build up the largest and most comprehensive segmentation dataset for training, by collecting over 22K 3D medical image scans from 72 segmentation datasets with careful standardization on both image scans and label space; (ii) for architecture design, we formulate a universal segmentation model, that can be prompted by inputting medical terminologies in text form. We present knowledge-enhanced representation learning on the combination of a large number of datasets; (iii) for model evaluation, we train a SAT-Pro with only 447M parameters, to segment 72 different segmentation datasets with text prompt, resulting in 497 classes. We have thoroughly evaluated the model from three aspects: averaged by body regions, averaged by classes, and average by datasets, demonstrating comparable performance to 72 specialist nnU-Nets, i.e., we train nnU-Net models on each dataset/subset, resulting in 72 nnU-Nets with around 2.2B parameters for the 72 datasets. We will release all the codes, and models in this work.
1 Introduction
Medical image segmentation involves identifying and delineating regions of interest (ROIs) like organs, lesions, and tissues in diverse medical images, which plays a crucial role in numerous clinical applications, such as disease diagnosis, treatment planning, and disease progression tracking [105, 113, 7, 79]. Traditionally, radiologists perform manual segmentation to measure volume, shape, and location in a slice-wise manner, which is both time-consuming and challenging to scale with the growing volume of medical data. Consequently, there is a pressing need for automated and robust medical image segmentation methods in clinical settings, to enhance efficiency and scalability.
Recent advancements in medical image analysis have been marked by a surge of deep learning. These developments have yielded a spectrum of segmentation models, each trained for specific tasks [71, 70, 5, 80, 7, 35, 69], often referred to as ‘specialist’ models. While these models demonstrate impressive segmentation capabilities, their major drawback lies in the narrow specialization. Designed and optimized for distinct ROIs and imaging modalities, these models [30, 32, 38, 75, 92] require distinct preprocessing methods for each dataset. As a result, they often fall short in diverse and dynamic clinical environments, where adaptability to new conditions and imaging techniques is essential.
Figure 1: Different from conventional specialist models (a) that develop specialized solution for each task, or recently proposed interactive universal segmentation models (b) relying on real-time human interventions, Segment Anything with Text (SAT) utilizes text prompts as queries and 3D volumes as inputs to perform a wide array of medical image segmentation tasks across different modalities, anatomies, and body regions (c). It can be easily applied to clinics or seamlessly integrated with any agent-based large language model.
There is a growing interest in developing universal models for medical image segmentation [68, 37], by adapting the pre-trained Segment Anything Model (SAM) [46] models from the computer vision community. However, while transferring to medical scenarios, these models trained on natural images suffer from fundamental limitations: (i) models typically perform 2D slice segmentation, that are later fused into 3D volumes through post-processing. This approach overlooks the critical contextual information inherent in 3D radiological imaging; (ii) models often require point or box inputs as prompts, thus is an interactive segmentation model, requiring considerable manual efforts for use in practice; (iii) models suffer from significant domain gaps, from image statistics to domain-specific medical knowledge, e.g., ‘liver is distinctive for positioning in the upper-right part of the abdominal cavity, filling almost the entire right hypochondrium and often extending into the left hypochondrium as far as the mammillary line’.
In this paper, we present the first knowledge-enhanced universal model for 3D medical volume segmentation, with language/text as prompt, termed as SAT. In practice, our model can effectively take 3D volume as visual inputs along with text prompts, to seamlessly tackle various medical image segmentation tasks, across modalities, anatomies, and body regions. As illustrated in Fig. 1, our proposed method distinguish from previous medical segmentation paradigms, can be seamlessly applied to clinics or integrated with any large language model. Specifically, we make the following contributions:
On dataset construction, we construct a knowledge tree based on multiple medical knowledge sources, encompassing thousands of anatomy concepts and definitions throughout the human body. On the visual side, we curate over 22K 3D medical image scans with 302K anatomical segmentation annotations, covering 497 categories from 72 publicly available medical segmentation datasets, termed as SAT-DS. To the best of our knowledge, SAT-DS denotes the largest and most comprehensive collection of public 3D medical segmentation datasets. To achieve such goal, we have invested significant efforts in standardizing datasets and unifying annotation labels, paving the way for universal segmentation task. For complete list of datasets and download links, we refer the reader to Table LABEL:tab:dataset_details and Table LABEL:tab:dataset_links.
Figure 2: Overview of our dataset for training universal segmentation model. Specifically, we collect datasets from a wide range of segmentation tasks across various modalities and anatomical regions of the human body, including the brain, head and neck, thorax, spine, abdomen, upper limb, lower limb, and pelvis.
On architecture design and training strategy, we build a universal medical segmentation model that leverages text prompts to guide segmentation, that enables flexible segmentation across a spectrum of medical imaging modalities. Specifically, we adopt knowledge-enhanced representation learning, leveraging textual anatomical knowledge and atlas segmentation of specific anatomical structure to train the text encoder. Through this training process, the visual features of these anatomical structures are aligned with their corresponding text descriptions in the latent space. Later, the text embeddings of anatomy/abnormality are treated as queries in a Transformer-based architecture, enabling the model to selectively extract task-specific features for precise segmentation of the queried target. To satisfy the requirements from different computational resources, we train two models of varying sizes, namely, SAT-Nano, and SAT-Pro.
On evaluation, we devise comprehensive metrics for universal medical segmentation across various aspects, including region-wise average, organ-wise average, and dataset-wise average. Through extensive experiments, we demonstrate that:
  • Building on the unprecedented dataset collection, SAT-Pro with only 447M parameters, can employ flexible text prompts for a wide range of downstream segmentation tasks, showing comparable performance to specialist nnU-Net models trained individually on each dataset, i.e., results from 72 separate nnU-Net models.
  • Benefiting from the standardized data processing and unified label system, SAT show excellent generalisation ability when performing zero-shot transfer to clinic data. It is worth emphasizing that SAT can flexibly satisfy the clinical needs of doctors from different departments, without any additional annotation and training.
  • Enhanced by multimodal medical domain knowledge, our text encoder provides superior guidance for universal medical segmentation on 3D inputs, surpassing the state-of-the-art language model tailored for medical tasks. For the first time, we demonstrate how domain knowledge would facilitate the segmentation task.
  • While enlarging model capacity, i.e., from SAT-Nano to SAT-Pro, would significantly improve the segmentation performance. This reveals the promising yet still under-exploited potential of a universal medical segmentation model.
  • The text-prompted feature distinguishes SAT from all previous universal medical segmentation models that require human interaction. We demonstrate how SAT can be directly integrated with large language models like GPT-4 [2], automatically segmenting the targets of interests without human intervention.
2 Results
The goal of our proposed SAT is to build an universal segmentation for 3D medical images, driven by text prompt. Such method should be capable of handling diverse segmentation tasks excellently in one model, without laborious specialization and training for each single dataset. In addition, the universality should make it adaptable to clinic procedures with minimal extra efforts, addressing a broad range of clinical needs; and the text-driven character should equip it with immense potential as a flexible toolkit, bridging between language and segmentation. Also, to meet different hardware requirements, we train two variants of SAT, namely SAT-Pro and SAT-Nano.
2.1 Main Results
Building an universal segmentation model poses several new challenges compared to specialist models, such as the heterogeneity of data, the long-tail distribution, the limitation from model capacity and so on. To this end, we have constructed a large-scale training set, namely SAT-DS dataset, covering a wide spectrum of medical image segmentation tasks, covering 497 anatomical targets on 8 regions and lesions of the human body, across 72 datasets. We train two universal segmentation models, namely, SAT-Pro and SAT-Nano, and compare them with the powerful specialist model, i.e., nnU-Nets trained on each dataset. To comprehensively understand such capability, we conduct the evaluations from the perspective of anatomical regions, classes, and datasets respectively. Note that, class-wise and region-wise evaluations require averaging results across different nnU-Nets. For example, the results on ‘brainstem’ in CT images is the macro average of nnU-Nets trained on ‘HAN Seg’, ‘PDDCA’ and ’SegRap2023 Task 1’, since all these datasets contains annotation for this class.
2.1.1 Region-wise Results.
Table 1 shows the segmentation performance of our methods and nnU-Nets on 8 regions of human body, including Brain, Head and Neck, Thorax, Abdomen, Pelvis, Spine, Upper Limb, and Lower Limb, as well as Lesion, in terms of DSC and NSD, respectively. Classes existing in multiple regions are specifically grouped as Whole Body. Notably, SAT-Pro consistently outperforms nnU-Nets in three regions: on Head and Neck, it surpasses nnU-Nets by 5.21 on DSC and 8.27 on NSD; On Upper Limb, it surpasses nnU-Nets by 2.8 on DSC and 9.28 on NSD; On Lower Limb, it leads to 1.53 on DSC and 2.48 on NSD. SAT-Nano rank last in almost all the regions, except for Spine, where it slightly exceeds SAT-Pro by 0.82 on DSC and 0.81 on NSD.
For comparison, we visualize the results in Fig. 3 (c). Overall, SAT-Pro shows comparable segmentation performance to the 72 nnU-Nets, that are specialized for each dataset, and outperforms SAT-Nano in most cases. In addition, it’s worth noticing that SAT-Pro not only performs better, more flexible, but also significantly smaller, approximately 1/5 of the assemble of nnU-Nets; While SAT-Nano is even smaller, remarkably only 1/20 of the assemble of nnU-Nets in size.
Figure 3: Comparison between SAT-Nano, SAT-Pro and nnU-Nets on different regions and lesions. (a) Dice score comparison; (b) NSD score comparison; (c) Model size comparison. SAT-Pro is approximately 5 times smaller than the total size of all the specialist nnU-Nets, while SAT-Nano is merely one twentieth the total size of all the specialist nnU-Nets.
Table 1: Segmentation results on different regions of human body and lesion. The best results are underlined. H&N: head and neck, LL: lower limb, UL: upper limb, WB: whole body.
Metric Method Brain H&N UL Thorax Spine Abdomen LL Pelvis WB Lesion
[…]
SAT-Pro 77.70 77.29 91.34 86.38 74.73 81.47 84.01 83.01 82.19 51.39
SAT-Nano 73.90 74.08 89.58 82.09 75.55 77.12 81.79 79.50 77.80 46.04
[…]
SAT-Pro 77.77 82.45 93.14 84.80 72.87 72.94 82.50 78.66 79.61 44.62
SAT-Nano 72.95 79.15 90.94 79.68 73.68 67.83 78.42 74.89 75.36 38.17
2.1.2 Class-wise Results.
For a more detailed evaluation, we show the results on all 497 classes in Fig. 4 and Fig. 5. Classes from the same region are clustered together and sorted from high to low according to the results of SAT-Pro. The scores of SAT-Pro, SAT-Nano, nnU-Nets on the same class are plotted in a vertical line. Statistics over all classes show that SAT-Pro outperforms SAT-Nano on most classes: 439/497 on DSC, and 442/497 on NSD; Meanwhile, SAT-Pro exceeds nnU-Nets on 133/497 classes on DSC, and 192/497 classes on NSD, including some important segmentation classes such as liver (+2.15 on DSC & +1.43 on NSD), pancreas (+0.83 & +2.76), lumbar vertebrae (+10.69 & +14.05). Averaging over all 497 classes, SAT-Pro achieves 78.73 on DSC, which is about 4.26% improvement over SAT-Nano (75.51); and 77.71 on NSD, about 5.31% improvement over SAT-Nano (73.76). Compared with nnU-Nets, SAT-Pro achieves 97.55% the performance of nnU-Nets on DSC and 99.32% the perfomance of nnU-Nets on NSD, demonstrating very competitive segmentation performance as an universal model. We report the detailed class-wise and dataset-wise results in Table 3, Table 4 and Table LABEL:tab:dataset_results.
Figure 4: DSC scores of SAT-Nano, SAT-Pro and nnU-Nets on each classes. Classes from the same region are clustered together and sorted from high to low according to the results of SAT-Pro.
Figure 5: NSD scores of SAT-Nano, SAT-Pro and nnU-Nets on each classes. Classes from the same region are clustered together and sorted from high to low according to the results of SAT-Pro.
2.2 Ablation Study
As will be illustrated in Sec. 6, to build an universal segmentation model driven by text prompts, we devise the visual encoder and decoder to process diverse medical images and perform segmentation; Meanwhile, domain knowledge is injected into the text encoder to provide precise prompts for the segmentation target, i.e., the encoding of terminology. In this section, we conduct experiments and discuss the effect of different visual backbones and domain knowledge in SAT. To save computational cost, all the experiment in this section are conduct on a subset of SAT-DS, termed ast SAT-DS-Nano. It includes 49 datasets, 13,303 images, 151,461 annotations and 429 classes.
2.2.1 Effect of Visual Backbone
Figure 6: Evaluations on different visual backbones on SAT-DS-Nano. ‘All’ denote the average scores over all the classes(n=429), including lesion classes. (a) DSC comparison; (b) NSD comparison. U-Net-CPT and U-Mamba-CPT perform very close, while both surpass SwinUNETR-CPT considerably.
Extensive studies have been done discussing the applicability of different visual backbones across segmentation tasks. However, it remains unexplored on which visual backbone suits the need of universal segmentation. In addition to the ConvNet-based U-Net, we consider two alternative backbones, namely, SwinUNETR [31] and U-Mamba [29] to medical segmentation. We configure three SAT-Nano variants with SwinUNETR(107M), U-Mamba(114M) and U-Net(110M) in comparable size: For SwinUNETR, we use the same hypeparameters as in official implementation [30]; For U-Mamba, we refer to the official implementation [67], reserve configuration of the U-Net in SAT-Nano, and only add mamba layers at the end of last 3 blocks of the U-Net encoder.
To avoid repeating multimodal knowledge injection on each visual backbones, we use MedCPT [42] for all these varients as text encoder, without loss of generality. Like the text encoder we pre-trained in Sec. 6.2, MedCPT is a text encoder trained on 255 million in-house user click logs from PubMed [91], and shows state-of-the-art performance on various biomedical language tasks. We expect it to possess domain knowledge and embed terminologies properly as text prompts. Thus, we denote these variants as U-NET-CPT, SwinUNETR-CPT and U-Mamba-CPT. We demonstrate the region-wise evaluation results in Fig. 6. U-Net-CPT outperforms U-Mamba-CPT slightly on both DSC (0.35) and NSD (0.22) scores averaged over all classes. However, both U-Net-CPT and U-Mamba-CPT exceed SwinUNETR-CPT by a significant margin. These observations confirms our choice of U-Net as the visual backbone of SAT. For more detailed comparison on each dataset and class, we refer the reader to Table LABEL:tab:visual_ab_dataset_results, Table 9 and Table 10.
2.2.2 Effect of Text Encoder
Figure 7: Evaluations on different text encoders on SAT-DS-Nano. ‘All’ denote the average scores over all the classes(n=429), including lesion classes. (a) DSC comparison; (b) NSD comparison. Prompted by MedCPT, SAT achieves better performance over BERT-base on average over all the classes. While our pre-trained text encoder outperforms MedCPT consistently across all the regions and lesion.
For an universal segmentation model with text as prompts, the neural representations of anatomical structures and lesions serve as the guidance for the segmentation model. And thus, the choice of text encoder should be carefully studied. In this section, we investigate the impact of domain knowledge on building a text encoder for medical universal segmentation task. Specifically, we train three SAT-Nano models with three representative text encoders: the text encoder pre-trained on our multimodal medical knowledge graph, as illustrated in Sec. 6.2; the state-of-the-art text encoder for various medical language tasks, i.e., MedCPT; And BERT-Base, a prevalent text encoder in natural language processing but not specifically finetuned on medical corpus. For all the variants, we use U-Net as visual backbone, and denote them as U-Net-Ours, U-Net-CPT, U-Net-BB.
Region-wise evaluation. As shown in Fig. 7, the performance of U-Net-CPT and U-Net-BB are close, with U-Net-CPT slightly outperfoms U-Net-BB on both DSC (+0.43) and NSD (+0.52) scores averaged over all classes, suggesting that finetuned on medical corpus, MedCPT might hold minor advantage. By contrast, U-Net-Ours surpasses U-Net-CPT consistently on all regions and lesions, leading notable margins on both DSC (+1.54) and NSD (+2.65) scores after average over all classes. This demonstrates the effectiveness of multimodal knowledge injection for training SAT. For more detailed comparison on each dataset and class, we refer the reader to Table LABEL:tab:text_ab_dataset_results, Table 6 and Table 7.
Concept-to-definition retrieval. To demonstrates advantage our proposed knowledge injection on capturing the critical features of medical terminologies, we evaluate the three text encoders on concept to definition retrieval task, with the human anatomy knowledge we collect, totaling 6,502 anatomy concepts and definition pairs. We find that the recall at 1 (R1) for BERT-Base is merely 0.08%, suggesting it can hardly understand these anatomy concepts or possess almost no domain knowledge; The R1 for MedCPT is 11.19%. Though this is a significant improvement over BERT-Base, MedCPT still struggles to distinguish these concepts; By contrast, our proposed text encoder get 99.18% R1, indicating the knowledge is successfully injected into the text embedding for each anatomy concept.
Knowledge injection helps segmentation of long-tail classes. The 429 classes in SAT-DS-Nano follow a typical long-tail distribution, as illustrated in Fig. 8 (a) and (b). The 10 classes with maximum annotations account for 12.75% of all the annotations in SAT-DS-Nano, denoted as ‘head’ classes; On contrast, the 150 classes with minimum annotations only accounts for 3.25%, even though they comprise 34.97% of 429 classes. We compare U-Net-Ours, U-Net-CPT and U-Net-BB on the ‘head’ classes, ‘tail’ classes and the rest ones (denoted as ‘middle’ classes). In Fig. 8 (c), we observe that prompted by our proposed text-encoder, U-Net-Ours achieves the best performance on all three type of classes. On ‘head’ classes, it outperforms the second best variant by 0.71 on DSC and 2.44 on NSD. While on ‘tail’ classes, the improvement is even more significant: 1.4 on DSC and 3.23 on NSD. Meanwhile, U-Net-CPT is better than U-Net-BB on ‘tail’ but slightly underperform on ‘head’. These results suggest our proposed text encoder is consistently better in a long-tail distribution, and the knowledge injection can effectively boost the segmentation performance on ‘tail’ classes.
Figure 8: The impact of domain knowledge on a long-tail distribution. (a) The annotation number of all the 429 classes in SAT-DS-Nano, sorted from highest to lowest. (b) The proportion of ‘head’, ‘middle’ and ‘tail’ in class number and annotation number. The DSC and NSD comparison on ‘head’, ‘middle’ and ‘tail’ classes.
2.3 Qualitative Results in Different Scenarios
Benefiting from the universality and text-driven features of SAT, it can be flexibly applied in clinic scenario. We verify such capability qualitatively on real clinic data. As opposed to previous universal segmentation model [46, 68, 14] that strictly relies on human interactions, SAT is built on knowledge enhanced text prompts, that can precisely understand the segmentation needs in text-form, regardless the user is a human or machine.
We utilize GPT-4 [2] to directly extract the anatomical targets of interests from real clinical report and prompt SAT to segment them on the clinical image, forming a fully automatic pipeline. In Fig. 9 and Fig. 10, we demonstrate the zero-shot performance of SAT-Pro on four cases randomly selected from clinical practice: abdominal MR examination, chest CT examination, abdominal CT examination, and lumbar spine CT examination. The basic information of patients, the report written by doctors, and targets of interests detected by GPT-4 are attached to each cases. As can be seen, the targets in reports are well detected by GPT-4 and commendably segmented by SAT-Pro: In Fig. 9, Liver, Pancreas, Kidneys, Gallbladder and Spleen are segmented in both abdominal CT and MR images; While in Fig. 10, segments of right lung, heart and bronchi are segmented in thoracic CT image, lumbar vertebrae L1 to L5 are segmented in spinal CT image. Note that in this evaluation, SAT-Pro is not fine-tuned on any data out of SAT-DS, despite the diversity of image and segmentation targets.
Figure 9: Zero-shot transfer to real clinical data. Top: A case from abdominal CT examination, no abnormal is observed from the CT image; Bottom: A case from abdominal MR examination, no abnormal is observed from the MR image; Basic information of patients, predefined targets of interests, and final reports written based on the segmentation results were appended to each cases.
Figure 10: Zero-shot transfer to real clinical data. Top: A case from chest CT examination, the patient is taking a follow-up examination after treatment for synovial sarcoma with pulmonary metastasis; Bottom: A case from lumbar vertebrae CT examination, some abnormalities are observed on the image; Basic information of patients, predefined targets of interests, and final reports written based on the segmentation results were appended to each cases.
3 Discussion
The heterogeneity of medical images has long called for versatile, effective and flexible universal medical segmentation model to meet clinical needs. To this end, we proposed SAT, a universal medical segmentation model trained on unprecedented dataset collection, and driven by knowledge enhanced text prompt. In this section, we discuss the performance, potential and limitations of the proposed SAT, based on our experiment results in Sec. 2.
An universal model is worth 72 specialist models. Through comprehensive experiments in both region-wise and class-wise evaluations, SAT-Pro has demonstrated comparable results to an assemble of 72 nnU-Nets specialized and trained on each dataset, and even surpass them on several regions and classes. These promising results reveal universal segmentation model as a much more flexible method qualified for wide range of tasks. Even more impressive, SAT-Pro is of much smaller size, implying a more cost-effective alternative to training and deploying a series of specialist models.
Scaling models lead to better performance. Increasing model capacity has been demonstrated to effectively enhance model’s performance in large language models [110, 99], yet such scaling law remains to be verified in segmentation tasks. In this work, we build SAT-Nano (110M) and SAT-Pro (447M). In both region-wise and class-wise evaluations, SAT-Pro shows clear performance boost over SAT-Nano, outperforming the latter on most region and classes. This implies that scaling law is also applicable to universal medical segmentation, probably a promising way to continuously improve the performance.
Domain knowledge can boost segmentation performance. The impact of domain knowledge, especially multi-modal knowledge on human anatomy remains under-explored for segmentation task. In this work, we propose a multi-modal knowledge graph on human anatomy, and via knowledge injection, build a text encoder prompting the segmentation task. In Sec. 2.2.2, we demonstrate how it can enhance the segmentation performance, especially on these ‘tail’ classes. As we expect the segmentation targets in human body to continuously grow in the future, it will raise higher demand on generating informative and distinguishable text prompts.
SAT can be easily transferred to clinics. Clinic utility is crucial for segmentation methods. In Sec. 2.3, we show that SAT-Pro can be applied directly to real clinical data out of the scope of SAT-DS without extra annotation and fine-tuning. Despite the diversity of image and segmentation targets from different clinic scenes, SAT can flexibly handle them with just one model. This could save laborious efforts to collect, annotate and train a bag of specialist models. More importantly, SAT can be utilized with text prompts, saving the time for per-sample human intervention, thus functioning automatically with predefined targets, reducing the cost in real use. Furthermore, as we have normalize all the image with unified voxel spacing, the successful application of SAT in clinic would pave the way for estimating the spacial size of anatomical targets directly, providing reference for analysis, diagnosis and treatment.
SAT can be a powerful agent for large language models. In Sec. 2.3, we show that SAT can be applied to segment targets extracted by GPT-4 from clinical report, providing explainable grounded report for the patients. This demonstrate the potential of SAT as a powerful agent, seamlessly integrated with any large language models, with text prompts bridging them. As grounding is an indispensable procedure in building generalist medical artificial intelligence [77], we expect the potential and significance of SAT to extend beyond a simple clinical tools.
Limitations. As an exploratory work in this field, though demonstrate impressive and promising results, we find certain limitation of SAT:
  • Firstly, the performance of SAT-Pro still lags behind nnU-Net in some region including Brain, Spine and Abdomen, and many classes especially the lesions. This implies SAT, though flexible and versatile, can not fully replace specialist models, especially in scenarios where absolute performance is critical. However, we evaluate SAT-Pro on train set, derive its averaged DSC score of all class as 83.68. We suspect its performance is still unsaturated, and continuous scaling-up universal segmentation model may close the gap and even outperforms the counterparts specialist models.
  • Secondly, SAT currently only supports text as prompts, thus not intended for scenes with human interaction requirements. For example, human annotation or refinement on the segmentation results.
  • Thirdly, the distribution of SAT-DS is unbalanced, leading to the limitation on capability of SAT. For instance, SAT-DS contains 189 classes for MR images, while 137 exists in ‘Brain’ region. Comprehensive annotation for MR images still lags behind CT images in the scope of public datasets.
  • Fourthly, because of the imbalance between different segmentation datasets, the long-tail distribution in assembled dataset collection remain challenging for building an universal segmentation method. As demonstrated in Fig. 8 (c), the performance gap between ‘head’, ‘middle’ and ‘tail’ classes is still significant.
4 Related work
4.1 Specialist Medical Image Segmentation
Recent years have witnessed significant advancements in the filed of medical image segmentation for diverse clinical applications [84, 9, 69, 97, 5, 8, 71, 43, 70, 35, 66, 106]. Since the success of convolutional neural networks (CNNs), U-Net [92, 16, 38] have shown dominant performance for medical image segmentation, and the variants [75, 21, 57, 115, 125, 94, 119] are further developed to dive into specific clinical tasks. Meanwhile, the emerging Transformers [32, 12, 30, 112, 117, 14, 122] step in this field and obtain promising results. This stream of specialist methods typically employ a customly-designed solution and a supervised learning framework. Despite the success in specific segmentation tasks, it requires to be trained from scratch on a new segmentation task, posing limitation on the generalisability. More importantly, the data scarcity for specific tasks often significantly hamper the application in diverse and practical clinical scenarios. In this paper, we present a generalist model that covers a wide spectrum of medical image segmentation tasks.
4.2 Generalized Medical Image Segmentation
In the Medical Segmentation Decathlon [5], nnU-Net [38] takes a promising step towards generalized medical image segmentation. It offers an automatic pipeline to define the optimal network architecture and data pre-processing for each segmentation task. However, the issues of specialist model remain. It has to involve substantial overhead by training from scratch on each task, and the intrinsic correspondence among segmentation tasks is neglected. To effectively benefit a multitude of clinical segmentation tasks, advanced learning schemes have been proposed for generalized medical image segmentation, including multi-task learning [83, 120, 26, 24], transfer learning [13, 19, 104], and continual learning [118, 41, 124]. Multi-task learning incorporates multiple imaging modalities, e.g.,multi-parametric MR and multi-phase CT [83, 120], or anatomies, e.g.,multi-organ segmentation [24]. In contrast, transfer learning achieves generalized medical image segmentation by transferring segmentation model from one task to another, e.g.,from CT to MR image segmentation and vice versa [13]. Continual learning incrementally involve novel tasks into the segmentation model and avoid forgetting the old ones at the same time [41]. However, these methods require either a limited pre-defined set of tasks or substantial tuning on novel tasks, and thus can hardly scale up. The proposed method exploits textual prompt learning, i.e.,the term of the organs or lesions, to smoothly steer the segmentation model to new tasks without costly training or tuning.
4.3 Universal Medical Image Segmentation
Universal medical image segmentation paves a way to superior generality and scalability by covering a wide range of organs and lesions across various medical imaging modalities and anatomies [37]. The universal model is usually driven by visual or textual prompts to adapt to various segmentation tasks. Since the significant success of SAM [46] in computer vision, several approaches follow the interactive framework and leverage box and/or point prompts to finetune the model on medical images [68, 37, 15, 102, 27]. Other methods also exploit image examples as visual prompts [11] or integrate visual prompts with textual prompts [22]. Despite showing promising results, these semi-automated approaches rely on carefully curated visual prompts, that may face critical challenges in clinical practice. Meanwhile, the typical 2D slice-wise segmentation in most visual prompting methods [68, 37, 15] inevitably leads to suboptimal context encoding for 3D medical imaging [27]. In contrast, pure textual prompts save the effort for burdensome interaction, especially when scaling up to tens of thousands of medical images. Several works have made successful attempts to adapt textual or learnable prompts for abdominal organ and lesion segmentation [61, 114], encouraging us to take a further step towards universal medical image segmentation. However, these studies still require a pre-defined set of categories and thus suffer from the shortage of medical domain knowledge, leading to limited generality across diverse medical segmentation tasks. To bridge this gap, we explicitly leverage multimodal medical knowledge composed of anatomy atlas and definition via visual language pre-training for universal medical image segmentation.
4.4 Knowledge-enhanced Representation Learning in Medical Image Analysis
As accurate medical image analysis relies on expert domain knowledge, several efforts have attempted to involve well-established medical knowledge to strengthen deep learning methods [111, 108]. These methods can be divided into two categories, namely, model-based and learning-based methods. Specifically, the model-based methods design the network architecture based on the clinical knowledge prior [55, 28] or mimicking the clinical practice [18, 23]. The learning-based methods exploit a multi-task learning scheme that takes medical knowledge as either inputs or additional supervision [107, 123]. Additionally, in the recent literature, [52, 89, 42] train or fine-tune language models with biomedical corpus [91, 10], boosting model performance over a wide range of biomedical language tasks; while [109, 116, 54, 121] adopt knowledge-enhanced visual-language representation learning and demonstrate superior performance in downstream classification tasks. Our method aligns with the representation learning methods. While all existing methods mainly focus on incorporation with domain knowledge in text form, we explore multimodal medical knowledge alignment for both text and anatomical atlas, facilitating fine-grained visual-language knowledge injection while training universal segmentation model.
5 Dataset
Toward building a universal segmentation model for medical imaging, with text as prompts, we collect two types of data in this work: medical domain knowledge to train the text encoder, and medical segmentation data, which includes both images and segmentation masks. In this section, we will describe them sequentially.
5.1 Domain Knowledge
To acquire textual knowledge, we mainly exploit two sources of domain knowledge: the Unified Medical Language System (UMLS) [^1] https://www.nlm.nih.gov/research/umls, a comprehensive medical knowledge graph consisting of concept definitions and their interrelations; and search engines, which are prompted to organize knowledge into a graph of the same format, specifically refined for the human anatomy corpus. Regarding visual knowledge, we have compiled 72 medical segmentation datasets, creating an atlas that covers over 497 anatomical structures of the human body. All sources of knowledge are systematically categorized, encompassing anatomies and concepts from all regions of the human body. Two examples of these sources are illustrated in Fig. 11 (a) and (b).
Figure 11: The medical knowledge used in the visual-language pretraining. (a) Segmentation datasets provide an atlas for extensive anatomy concepts. In this example, atlas segmentation is marked with different colors. (b) Knowledge generated from UMLS and search engines encompasses a broad array of concept-definition pairs and extensive relationships. (c) By integrating all collected knowledge sources, a medical knowledge tree is constructed. All definitions are partially displayed for conciseness. Definition and relation denoted with * are derived from the search engine, otherwise from UMLS.
UMLS. Unified Medical Language System (UMLS) [10] is a knowledge source of biomedical vocabulary developed by the US National Library of Medicine [^2] https://www.nlm.nih.gov. It integrates a wide range of concepts from more than 60 families of biomedical vocabularies, each equipped with a Concept Unique Identifier (CUI) and definition. It also contains the relations among these concepts. Following [116], we extract 229,435 biomedical terminologies and definitions, as well as 1,048,575 relationship triplets, composing a knowledge graph on these terminologies.
Search engine generared knowledge. Although UMLS is widely acknowledged and adopted as a general medical knowledge corpus [109, 116, 54, 121], it lacks a fine-grained description of anatomy concepts critical for segmentation tasks. For example, for ‘liver’, the definition is ‘A large lobed glandular organ in the abdomen of vertebrates that is responsible for detoxification, metabolism, synthesis, and storage of various substances.’, which erases the morphological features and focuses on functionality. Meanwhile, more comprehensive knowledge on human anatomy may be scattered across various authoritative websites online, e.g., Wikipedia [^3] https://www.wikipedia.org, ScienceDirect [^4] https://www.sciencedirect.com/browse/journals-and-books?subject=medicine-and-dentistry, etc.. To harvest such knowledge, we select 6,502 anatomy concepts, and prompt an search engine  [^5] https://you.com to retrieve and summarize definitions for them. We use the following prompt template: {mdframed}[backgroundcolor=gray!10, linewidth=0pt] Definition of xxx. Include the location, shape, appearance, structure and spatial relations to other anatomy structures. No need to include functionality. End with ‘END’. For illustration, the search engine referred to authority websites including Columbiasurgery [^6] https://columbiasurgery.org/liver/liver-and-its-functions, Hopkins Medicine [^7] https://www.hopkinsmedicine.org/health/conditions-and-diseases/liver-anatomy-and-functions and summerized definition for ‘liver’ as: ‘A large organ found in the upper right quadrant of the abdomen, it stands as the largest gland within the human body, with a weight of about 1.5 kilograms. This structure exhibits a reddish-brown hue and is cone or wedge-shaped … …’. While constructing the knowledge graph, we also adopt GPT4 [2] to extract 38,344 relations between anatomical structures in the generated information-dense definitions with the following prompt: {mdframed}[backgroundcolor=gray!10, linewidth=0pt] This is the description of xxx. Please help me find its relations with other anatomical structures in radiological images. Summarize them with the template: Relation: xxx (relational preposition), Anatomical structure: xxx (name of another anatomical structure).
For example, “Relation: situated below, Anatomical structure: xxx”, “Relation: connected to (via xxx), Anatomical structure: xxx” … …
Visual Atlas. Segmentation datasets naturally provide visual features for anatomy concepts corresponding or complementary to the textual description, such as the texture, spatial location, shape and so on. Details on our collected segmentation datasets are described in Sec. 5.2. Here, we use them as a large-scale and diverse visual atlas library, and link the visual regions to corresponding concepts in textual knowledge graph, bridging the knowledge between visual and language modality.
In summary, by mixing these data, we construct a multimodal medical knowledge tree. As demonstrated in Fig. 11 (c), the concepts (including both anatomical structures and lesions) are linked via the relations and further extended with their definitions, containing their characteristics. Additionally, some are further mapped to the visual atlas, demonstrating their visual features that may hardly be described purely by text.
5.2 Segmentation Dataset
To train our segmentation model with the ability to handle segmentation tasks of different targets, across various modalities and anatomical regions, we collect and integrate 72 diverse publicly available medical segmentation datasets, totaling 22,186 scans including both CT and MRI and 302,033 segmentation annotations spanning 8 different regions of the human body: Brain, Head and Neck, Upper Limb, Thorax, Abdomen, Plevis, and Lower Limb. The dataset is termed as SAT-DS. More details are present in Table LABEL:tab:dataset_details and Table LABEL:tab:dataset_links. Note that, some public datasets are not mutually exclusive, e.g., KITS23 and KITS21 [35], we thus only collect the latest version, to avoid redundancy and potential data leakage in train-test split.
Before mixing these datasets for training, two challenges remain: (i) the anatomical targets from each dataset must be integrated into a unified annotation system. The clinic demands beneath each dataset collection might be different, resulting in different annotation standards and granularity. Meanwhile, since most datasets are annotated for training specialist models like nnU-Net [38], precise and consistent terminology or expression for anatomical targets is often ignored. Therefore, a unified label system is demanded to avoid potential contradictions when training on mixed datasets. (ii) some critical image statistics, such as intensity distribution and voxel spacing vary from dataset to dataset, hindering the model from learning consistent image representations across datasets. In the following, we present details for dataset integration and pre-processing, and how we address the abovementioned challenges.
Figure 12: Statistics of  SAT-DS across different anatomical regions. (a) Annotation num of some representative classes in each anatomical region; (b) Number of classes in each anatomical region; (c) Number of annotations in each anatomical region. LC/HC: laryngeal/hypopharyngeal cancer, WHM: white matter hyperintensities, PV&SV: portal vein and splenic vein, IVC: inferior vena cava, Thoracic V: thoracic vertebrae, Cervical V: cervical vertebrae, Lumbar V: lumbar vertebrae.
[…]
Dataset Pre-processing. As properties of each dataset may greatly impact the training of the segmentation network [38], such as intensity distribution and voxel spacing, we deliberately apply some normalization procedures to all the datasets to ensure uniformity and compatibility between them. Firstly, all the images are reoriented to specific axcodes, respaced to a voxel size of $1\times 1\times 3\leavevmode\nobreak\ mm^{2}$ and cropped to non-zero region. Secondly, we apply different intensity normalization strategies to CT images and MRI images, specifically, for CT images, intensity values are truncated to $[-500,1000]$ and applied z-score normalization. For MRI images, intensity values are clipped by $0.5\%$ and $99.5\%$ of the image, and then z-score normalized. During training, we randomly crop the image patch with a fixed size of $288\times 288\times 96$. Random zoom-in, zoom-out, and intensity scaling are applied for data augmentation.
Dataset Statistics. After integrating datasets, we derive a segmentation data collection that covers 497 segmentation classes, far outpacing each single dataset in both diversity and scale. We divide human body into eight regions, and classify each class into them manually. Fig. 12 (b) and (c) show the distribution of classes and annotations across different human body regions. We further show the distribution of some example classes in each region in Fig. 12 (a). The extensive range of categories and regions lays the foundation for the SAT’s wide application scenarios. The data collection is more than fourth times the size of the largest dataset (BraTS2023-GLI) in terms of volume number, and nearly triple the most comprehensive dataset (DAP Atlas) in terms of the class number.
Discussion. In the process of building SAT-DS, we merge a wide range of segmentation tasks, and establishing a unified label system by using natural language/text. Generally speaking, there are three advantages to do this: (i) natural language is powerful and discriminative, that enables better differentiation of the medical terminologies in the language embedding space; (ii) as shown in previous work [109, 116, 54, 121], knowledge-enhanced representation learning for the text encoder demonstrates promising performance, allowing to learn the implicit or explicit relationships between these segmentation targets. For example, segmenting a specific lobe of the liver requires the exact segmentation of liver as an organ in abdomen cavity, and shall be facilitated by referring to other parts of the liver. Therefore, establishing such connections via systematic medical knowledge shall be beneficial. (iii) text prompt can be given automatically without any human intervention, for instance, from large language model. This would pave the way for building a universal segmentation model that can be flexibly integrated into foundation models for generalist medical artificial intelligence, as a powerful grounding tool.
6 Method
Towards building up our universal segmentation model by text prompt, we consider two main stages, namely, multimodal knowledge injection and universal segmentation training. In the following, we start by defining the problem scenario (Sec. 6.1). Then, we show how to structure multimodal medical knowledge and inject it into a text encoder (Sec. 6.2). Lastly, we employ the text encoder to guide our universal segmentation model training on SAT-DS dataset (Sec. 6.3).
6.1 Problem Formulation
Assuming we have a segmentation dataset collection, i.e., $\mathcal{D}=\{(x_{1},y_{1};T_{1}),...,(x_{K},y_{K};T_{K})\}$, where $x_{i}\in\mathbb{R}^{H\times W\times D\times C}$ denotes the image scan, $y_{i}\in\mathbb{R}^{H\times W\times D\times M}$ is the binary segmentation annotations of the anatomical targets in the image, and $T_{i}=\{t_{1},t_{2},...,t_{M}\}$ denote the corresponding medical terminology set. The universal segmentation task can be formulated as:
$\hat{y}_{i}=\Phi_{\text{SAT}}(\Phi_{\text{visual}}(x_{i}),\Phi_{\text{text}}(T_{i})),$ (1)
where $\Phi_{\text{visual}}$ is a visual encoder, $\Phi_{\text{text}}$ is a text encoder, $\Phi_{\text{SAT}}$ is a universal segmentation model and ideally, $x_{i}$ can be an image scan from any modality of the anatomical region, $T_{i}$ can contain any text-wise medical terminology of interest.
6.2 Multimodal Knowledge Injection
Here, we aim to inject rich multimodal medical knowledge into the visual and text encoders. The section starts from the procedure for structuring the multimodal medical knowledge data, and further presents details to use them for visual-language pre-training.
Multimodal Domain Knowledge Construction. As shown in Fig. 13 (a), the data from UMLS, search engine, and segmentation datasets can be aggregated into two formats:
  • Textual medical concept pair. For text-only knowledge, each concept $t_{i}$ is associated with a definition $p_{i}$, constructing pairs of text $(t_{i};p_{i})$. We also derive a knowledge graph that connects the medical concepts through abundant triplet relationships $(t_{i},r_{ij},t_{j})$. This graph can be alternatively seen as a specialized text pair, $(t_{i}+r_{ij};t_{j})$ or $(t_{i};r_{ij}+t_{j})$, where ‘$+$’ refers to string concatenation. In this way, we can thus unify the two kinds of textual knowledge.
  • Visual medical concept pair. To align with the segmentation task, we gather pairs consisting of a concept (can be either an anatomical structure or lesion) and its image atlas. Note that, multiple pairs could be extracted from a single image. These pairs share a similar format to the segmentation data, denoted as $(x_{i},y_{i};t_{i})$, where $x_{i}$ and $t_{i}$ are consistent with their definition in Sec. 6.1 and $y_{i}\in\mathbb{R}^{H\times W\times D\times 1}$ is a binary segmentation mask for $t_{i}$.
In summary, all the knowledge can either be represented as pure text description, e.g., $t_{i}$, $p_{i}$, $t_{i}+r_{ij}$, $r_{ij}+t_{j}$, or atlas segmentation $(x_{i},y_{i})$, and paired further.
Figure 13: Overview of SAT. (a) We inject multi-modal medical knowledge into knowledge encoders via contrastive learning. The knowledge is in different formats: atlas segmentation, concepts(terminologies), definitions, and relationships between concepts. We devise visual and text encoder to embed them; (b) We train a universal segmentation network based on the text prompts from the pre-trained text encoder. It is capable of segmenting a wide range of targets for image scans from different modalities and anatomical regions.
Visual-Language Pre-training. As shown in Fig. 13 (a), for pure text description, we encode them with a BERT [45] pre-trained on PubMed abstract [52]:
$z=\Phi_{\text{text}}(\textbf{t}),\text{ }\textbf{t}\in[t_{i},p_{i},t_{i}+r_{ij},r_{ij}+t_{j}],\text{ }z\in\mathbb{R}^{d},$ (2)
where $d$ refers to the feature dimension. For visual concepts, we adopt the visual encoder $\Phi_{\text{visual}}$. Given the excellent robustness and performance of U-Net [38, 92], we apply a standard 3D U-Net encoder to extract multi-scale image embeddings:
$V_{i}=\{v_{i1},v_{i2},...,v_{iS}\}=\Phi_{\text{visual}}(x_{i}),\text{ }v_{is}\in\mathbb{R}^{H_{s}\times W_{s}\times D_{s}\times d},$ (3)
where $V_{i}$ is the multi-scale feature maps from U-Net encoder layers, and $H_{s},W_{s},D_{s},d_{s}$ are the spatial resolutions and channel width at different layers. We further average ROI pooling on these feature maps respectively, based on the down-sampled segmentation mask fitting resolutions at different layers. The concatenated pooled features can thus be treated as representation of the anatomical target on this image, containing multi-scale visual clues for it.
$z=\mathcal{F}_{\text{pooling}}(\Phi_{\text{visual}}(x_{i});y_{i}),\text{ }z\in\mathbb{R}^{d}.$ (4)
We train the visual encoder and text encoder by maximizing the similarities between all positive knowledge pairs, linked by text prompts (medical terminology names), as shown in Fig. 13. Specifically, given $(x_{i},y_{i};t_{i}),(t_{i};p_{i}),(t_{i}+r_{ij};t_{j}),(t_{i};r_{ij}+t_{j})$, for simplicity, we denote all the encoded features as $z$, regardless of their modality format. For a batch of $N$ pairs $\{(z_{1},z_{1}^{\prime}),...(z_{N},z_{N}^{\prime})\}$, we have:
$\mathcal{L}_{\text{knowledge}}=-\frac{1}{N}\sum^{N}_{i=1}(\log\frac{\exp(z_{i}\cdot z_{i}^{\prime}/\tau)}{\sum^{N}_{k=1}\mathbbm{1}_{i\neq k}\exp(z_{i}\cdot z_{k}^{\prime}/\tau)}+\log\frac{\exp(z_{i}\cdot z_{i}^{\prime}/\tau)}{\sum^{N}_{k=1}\mathbbm{1}_{i\neq k}\exp(z_{k}\cdot z_{i}^{\prime}/\tau)}),$ (5)
which is a typical contrastive learning pipeline, with $\tau=0.07$ as temperature.
Discussion. In general, by maximizing the similarities between positive textual and visual feature pairs, we force the text encoders to construct neural representations for medical concepts from two aspects: (i) through the well-established knowledge graph in text form, the text encoder enables to encode relationships between concepts in the latent space; (ii) the model captures the characteristics of anatomical structures and lesions via both visual atlas segmentations and detailed definitions. Therefore, in contrast to the one-hot labeling that treats each anatomical target as of being independent, such continuous neural representation shall provide more helpful guidance for the segmentation task.
6.3 Universal Segmentation Training
With the pre-trained visual and text encoder, we now continue the procedure for building the universal segmentation model with text as prompts. Fig. 13 (b) demonstrates the overall framework. Specifically, apart from the pre-trained visual and text encoders, the segmentation model consists of three more components: a visual decoder $\Phi_{\text{dec}}$, a query decoder $\Phi_{\text{query}}$, and a mask generator.
Although a sample in the segmentation dataset collection $(x_{i},y_{i};T_{i})$ may contain multiple annotations, i.e., $T_{i}=\{t_{1},t_{2},...,t_{M}\}$, for simplicity, we describe the segmentation procedure for one structure ($t_{i}$).
Terminology Encoding. We use the pre-trained text encoder to generate neural embedding for one anatomical target as a text prompt for segmentation:
$z_{i}=\Phi_{\text{text}}(t_{i}),\text{ }z_{i}\in\mathbb{R}^{d}.$ (6)
Note that, after pre-training the text encoder with domain knowledge injection, $z_{i}$ should contain both the textual background information and visual information from atlas samples.
Visual Encoding. We adopt the pre-trained visual encoder as explained in Sec. 6.2, Equa. 3, and continue training it.
Visual Decoding. In the visual decoder, the feature maps from the encoder are gradually upsampled with skip connections, effectively following the U-Net architecture [38, 92], ending up with per-pixel dense feature:
$u_{i}=\Phi_{\text{dec}}(V_{i}),\text{ }u_{i}\in\mathbb{R}^{H\times W\times D\times d^{\prime}},$ (7)
where $d^{\prime}$ is the dimension for the per-pixel dense feature after recovering to the original resolution.
Query Decoder. Although a general representation of the anatomical target is derived from the pre-trained text encoder with a text prompt, visual variations may still exist from patient to patient, we thus insert a transformer-based query module to further enhance the text prompts with visual clues. In practice, this module consists of 6 standard transformer decoders [101], that treats text embedding as query, and the pooled multi-scale visual features from the U-Net encoder as key, values, formulated as:
$q_{i}=\Phi_{\text{query}}(V_{i},z_{i}),\text{ }q_{i}\in\mathbb{R}^{d}.$ (8)
Where $z_{i}$ is consistent with $z$ in Equa. 4. Therefore $q_{i}$ can be seen as an adapted representation of the anatomical target in specific image scan $x_{i}$.
Mask Generator. By conducting pixel-wise dot-product between the representation of the anatomical target and the fine-grained per-pixel embedding, we can acquire a per-pixel segmentation:
$\hat{y}_{i}=g(q_{i})\cdot u_{i},\text{ }\hat{y}_{i}\in\mathbb{R}^{H\times W\times D},$ (9)
where $g(\cdot)$ is a feed-forward layer projecting $q_{i}$ to a consistent dimension with the dense feature map $u_{i}$.
Training Objective. Following [38], we adopt a loss function as the sum of binary cross-entropy loss and dice loss. For a sample with $M$ classes and $C$ voxels, we denote $p_{\text{c},\text{m}}$ and $s_{\text{c,m}}$ the prediction and ground-truth for $c$-th pixel respectively on class $m$, the loss is:
$\mathcal{L}=\underbrace{-\frac{1}{M}\sum_{\text{m=1}}^{M}\frac{1}{C}\sum_{\text{c=1}}^{C}p_{\text{c,m}}\cdot\log s_{\text{c,m}}}_{\text{Binary Cross Entropy Loss}}+(\underbrace{1-\frac{2\sum_{\text{i=1}}^{M}\sum_{\text{c=1}}^{C}p_{\text{c,m}}\cdot s_{\text{c,m}}}{\sum_{\text{m=1}}^{M}\sum_{\text{c=1}}^{C}p_{\text{c,m}}^{2}+\sum_{\text{m=1}}^{M}\sum_{\text{c=1}}^{C}s_{\text{c,m}}^{2}}}_{\text{Dice Loss}})$ (10)
7 Implement Details
We encounter several challenges in training on the combined large number of heterogeneous medical datasets in 3D format. In this section, we provide details on the adopted training strategies.
Progressive Multimodal Knowledge Injection. For the stability of training, we implement the multimodal knowledge injection procedure introduced in Sec. 6.2 progressively. At the beginning, the text encoder is pretrained on the text knowledge via contrastive learning, i.e., the textual medical concept pairs. The maximal sequence length after tokenization is 256, for even longer text input, we apply random truncation to fully exploit the knowledge in the long text. After convergence, to align text representations and visual features, we apply contrastive learning between the finetuned text encoder and the visual encoder on the visual medical concept pair.
Pre-processing Segmentation Dataset. We take voxel spacing as $1\times 1\times 3\ mm^{2}$ and patch size $288\times 288\times 96$, based on two empirical considerations: (i) when mixing the various datasets, scans with a wide range of voxel spacings ought to be normalized before processing in the same convolutional network. While resampling to larger voxel spacing may lose information, smaller voxel spacing will generate artifacts; (ii) a larger receptive field generally ensures better segmentation performance for most targets, however, increasing patch size will result in higher computational cost. We strike a balance between them based on the computational resources in use.
Balancing Segmentation Datasets. To balance between all datasets, we set the sampling strategy for each scan based on two intuitions respectively: (i) training case number varies significantly from dataset to dataset, which should be alleviated. We follow  [100] and set the sampling weight of all scans in a dataset as the inverse proportion to $\sqrt{N}$, $N$ is the number of training cases in the dataset; (ii) scans with larger annotation areas or more annotated classes should be sampled more as they are often harder to learn. Thus, we repeat such scans for $R$ times in the sampling pool, where $R=\frac{S_{roi}}{288\times 288\times 96}\times\frac{M}{32}$. $S_{roi}$ is size of the annotated area in a image scan, namely, its foreground area; $288\times 288\times 96$ is the patch size to crop; $c$ is the number of annotated classes on it; and 32 is the maximal number of text prompts in a batch.
Balancing Segmentation Classes. While cropping the image scan, it’s a common practice [38] to over-sample foreground crops, i.e., crops containing at least one segmentation target. However, weighting these crops evenly may ignore the unbalanced spatial distribution of segmentation targets. For example, in large scans with numerous annotations, tiny targets are harder to sample and thus may be ignored by the model. Thus, in foreground oversampling, we give more weight on regions with more segmentation targets.
Difference between SAT-Pro and SAT-Nano. We devise two variants of SAT with different model sizes in visual backbone. For SAT-Nano, we adopt a U-Net with 6 blocks in-depth, each block has 2 convolutional layers and $3\times 3$ size each kernel. The channel widths for each stage are $[64,64,128,256,512,768]$; SAT-Pro share the same architecture with SAT-Nano, except that each block consists of 3 convolutional layers, and the channel widths are doubled to $[128,128,256,512,1024,1536]$.
Other Hyperparameter Details. The query decoder is a 6-layer standard transformer decoder with 8 heads in each attention module. Feature dimensions of a text prompt $d=768$ and per-pixel embedding $d^{\prime}=64$. To unify the features from visual backbone variants with varying channel widths, they are projected to $d=768$ with different feed-forward layers, and input as key and value to query decoder. For images with multiple segmentation targets, we set the maximal text prompts sampled in a batch of up to 32. A combination of cross-entropy loss and dice loss is applied as supervision at training time. We use AdamW [63] as the optimizer with cosine annealing schedule, maximal $lr=1\times 10^{-4}$, 10000 steps for warm-up. The SAT-Nano is trained on 8 NVIDIA A100 GPUs with 80GB memory, using maximal batch size 2; while the SAT-Pro is trained on 16 NVIDIA A100 GPUs with 80GB memory, using maximal batch size 1.
8 Experiment Settings
8.1 Evaluation Datasets
To strike a balance between extensive experiments and computational cost, we utilize two collection of datasets:
  • SAT-DS. As describe in Sec. 5.2, this contains all the 72 datasets, 497 classes from all human body regions, 22,186 image scans and 302,033 segmentation annotations.
  • SAT-DS-Nano. A subset of SAT-DS, including only 49 datasets, 13,303 images and 151,461 annotations. Note that SAT-DS-Nano also covers 429 classes from all human body regions, adequate to evaluate universal segmentation task.
The detailed composition of SAT-DS and SAT-DS-Nano can be found in Table LABEL:tab:dataset_details. As there is no existing benchmark for evaluating the universal segmentation model, we randomly split each dataset into 80% for training and 20% for testing: (i) datasets may share the same images but with different classes. For example, Couinaud Liver provides fine-grained liver segmentation on a subset of MSD Hepatic Vessel. We carefully split the Couinaud Liver to make sure the test set will not be leaked in the train set of MSD Hepatic Vessel; (ii) scans of the same patient but different modalities are treated as different samples during training and evaluation. For example, MSD Prostate contains T2 and ADC scans of each patient. However, they share the same structure on the image. To avoid potential data leaking, we carefully split such datasets by patient id. Note that when involving segmentation datasets in the visual-language pretraining, we only use the training data to avoid potential data leaking. For datasets involved in SAT-DS-Nano, we keep their splits the same as in SAT-DS. The download link for each datasets can be found in Table LABEL:tab:dataset_links, and we will release our dataset splits to the research community for model comparison.
8.2 Baseline
We take nnU-Net [38] as a strong baseline for comparison, which specializes in a series of configurations for each dataset. For a comprehensive evaluation, we train one nnU-Net model on each of the datasets. Note that, following [106], we split Totalsegmentator into 6 subsets and train 6 different nnU-Net models. Similarly, datasets such as CHAOS with both MRI and CT images are treated as two different datasets. When training nnU-Nets on each dataset, we adopt a multi-class segmentation setting and deliver the masks of all categories in this dataset at once. We derive the optimal network architecture and pre-processing pipeline with the default setting of nnU-Net. The detailed network design is shown in Table LABEL:tab:nnunet_config. In summary, we train an assemble of 72 nnU-Nets, that are customized on each dataset and serve as the competitive baseline for evaluation. We adopt the latest official implementation of nnU-Net v2888https://github.com/MIC-DKFZ/nnUNet/ in practice.
8.3 Evaluation Protocols
Given our goal is to develop a universal medical segmentation model, this provides opportunities to evaluate novel perspectives in addition to the traditional evaluation per dataset. Specifically, we conduct the evaluations from three dimensions:
  • Class-wise Evaluation. As SAT is capable of segmenting a wide range of anatomical targets across the human body, we merge the results from the same classes across datasets to indicate the performance on each anatomical target. Specifically, we follow macro-average method: for a class annotated in multiple datasets, we first calculate its average scores within each dataset, and then average them over all datasets. Note that, the same anatomical structures or lesions from different modalities are treated as different classes in this work, e.g., liver in both CT and MRI images.
  • Region-wise Evaluation. In general, anatomical structures from the same human body region are closely connected and more likely to be involved in diagnosis within the same hospital department. Here, we consider the region-wise evaluation: based on class-wise evaluation, we merge results from all classes in the same body region, as to indicate the general performance in this region. For classes existing in multiple regions, we classify them into ‘Whole Body’ category. In addition, we report results for lesions classes independently as a category ‘lesion’, instead of merging them into specific regions.
  • Dataset-wise Evaluation. Results of the classes within the same dataset are averaged to indicate the performance on this dataset. This is the same as the conventional evaluation protocol of specialist segmentation models trained on a single dataset, however, we believe this is less valuable for evaluating universal segmentation models.
8.4 Evaluation Metrics
We quantitatively evaluate the segmentation performance from the perspective of region and boundary metrics [72], e.g., Dice Similarity Coefficient (DSC) and Normalized Surface Distance (NSD) respectively.
DSC. Dice Similarity Coefficient (DSC) is a standard region-based metric for medical image segmentation evaluation. It measures the overlap between model’s prediction $P$ and ground truth $G$, formally defined as:
$DSC(P,G)=\frac{2|P\bigcap G|}{|P|+|G|}.$ (11)
NSD. Normalized Surface Distance (NSD) [78] is a boundary-based metric that measures the consistency at the boundary area of the model’s prediction $P$ and ground truth $G$, which is defined as:
$NSD(P,G)=\frac{|\partial P\bigcap B_{\partial G}|+|\partial G\bigcap B_{\partial P}|}{|\partial P|+|\partial G|},$ (12)
where $B_{\partial P}=\{x\in\mathbf{R}^{3}|\exists\hat{x}\in\partial P,||x-\hat{x}||\leq\tau\}$ and $B_{\partial G}=\{x\in\mathbf{R}^{3}|\exists\hat{x}\in\partial G,||x-\hat{x}||\leq\tau\}$ are the boundary area of model’s prediction and ground truth at a tolerance $\tau$, respectively. We set $\tau$ as 1 in the experiments.
9 Conclusion
In this paper, we promote the progress of medical universal segmentation with text as prompt and knowledge enhancement. We build up the largest and most comprehensive 3D medical segmentation datase, and the first multi-modal knowledge tree for human anatomy. In contrast to existing segmentation models that are limited in 2D input and box/point prompts, our model, SAT, enables text-driven universal segmentation on various radiologic modalities across 8 main human body parts, greatly reducing the usage difficulty of medical universal segmentation model in practice. The final solution we proposed, i.e., SAT-Pro, which only contains 447M paramters, while demonstrating comparable performance to 72 specialist nnU-Nets, with around 2.2B parameters trained individually on 72 datasets for segmenting 497 categories. We demonstrate its strong generalization capability to clinic data, as an off-the-shelf automatic segmentation method serving the doctor in analysis and diagnosis. Importantly, we show the potential of SAT as a powerful agent, paving the way to equip large language models with grounding capability.