### v3

# One Model to Rule them All: Towards Universal Segmentation for Medical Images with Text Prompts

Abstract
Abstract. In this study, we aim to build up a model that can Segment Anything in radiology scans, driven by Text prompts, termed as SAT. Our main contributions are three folds: (i) for dataset construction, we construct the first multi-modal knowledge tree on human anatomy, including 6502 anatomical terminologies; Then we build up the largest and most comprehensive segmentation dataset for training, by collecting over 22K 3D medical image scans from 72 segmentation datasets, across 497 classes, with careful standardization on both image scans and label space; (ii) for architecture design, we propose to inject medical knowledge into a text encoder via contrastive learning, and then formulate a universal segmentation model, that can be prompted by feeding in medical terminologies in text form; (iii) As a result, we have trained SAT-Nano (110M parameters) and SAT-Pro (447M parameters), demonstrating comparable performance to 72 specialist nnU-Nets trained on each dataset/subsets. We validate SAT as a foundational segmentation model, with better generalization ability on external (unseen) datasets, and can be further improved on specific tasks after fine-tuning adaptation. Comparing with interactive segmentation model, for example, MedSAM, segmentation model prompted by text enables superior performance, scalability and robustness. As a use case, we demonstrate that SAT can act as a powerful out-of-the-box agent for large language models, enabling visual grounding in clinical procedures such as report generation. All the data, codes, and models in this work have been released.
1 Introduction
Medical image segmentation involves identifying and delineating regions of interest (ROIs) like organs, lesions, and tissues in diverse medical images, which plays a crucial role in numerous clinical applications, such as disease diagnosis, treatment planning, and disease progression tracking [89, 94, 7, 65]. Traditionally, radiologists perform manual segmentation to measure volume, shape, and location in a slice-wise manner, which is both time-consuming and challenging to scale with the growing volume of medical data. Consequently, there is a pressing need for automated and robust medical image segmentation methods in clinical settings, to enhance efficiency and scalability.
Recent advancements in medical image analysis have been marked by a surge of deep learning. These developments have yielded a spectrum of segmentation models, each trained for specific tasks [57, 56, 5, 66, 7, 24, 55], often referred to as ‘specialist’ models. While these models demonstrate impressive segmentation capabilities, their major drawback lies in the narrow specialization. Designed and optimized for distinct ROIs and imaging modalities, these models [19, 21, 27, 61, 77] require distinct preprocessing methods for each dataset. As a result, they often fall short in diverse and dynamic clinical environments, where adaptability to new conditions and imaging techniques is essential.
Figure 1: In contrast to conventional specialist models (a) that develop specialized solution for each task, or recently proposed interactive universal segmentation models (b) relying on real-time human interventions, Segment Anything with Text (SAT) directly takes 3D volumes as inputs, and use text as prompts to perform a wide array of medical image segmentation tasks across different modalities, anatomies, and body regions (c). It can be easily applied to clinics or seamlessly integrated with any agent-based large language model.
There is a growing interest in developing universal models for medical image segmentation [54, 26], by adapting the pre-trained Segment Anything Model (SAM) [34] models from the computer vision community. However, while transferring to medical scenarios, these models trained on natural images suffer from fundamental limitations: (i) models typically perform 2D slice segmentation, that are later fused into 3D volumes through post-processing. This approach overlooks the critical contextual information inherent in 3D radiological imaging; (ii) models often require point or box inputs as prompts, thus is an interactive segmentation model, requiring considerable manual efforts for use in practice; (iii) models suffer from significant domain gaps, from image statistics to domain-specific medical knowledge.
In this paper, we present the first knowledge-enhanced universal model for 3D medical volume segmentation, with language/text as prompt, termed as SAT. In practice, our model can effectively take 3D volume as visual inputs along with text prompts, to seamlessly tackle various medical image segmentation tasks, across modalities, anatomies, and body regions. As illustrated in Fig. 1, our proposed method distinguishes from previous medical segmentation paradigms, that can be seamlessly applied to clinics or integrated with any large language model. Specifically, we make the following contributions:
On dataset construction, we construct a knowledge tree on anatomy concepts and definitions throughout the human body. On the visual side, we curate over 22K 3D medical image scans with 302K anatomical segmentation annotations, covering 497 categories from 72 publicly available medical segmentation datasets, termed as SAT-DS. To the best of our knowledge, SAT-DS denotes the largest and most comprehensive collection of public 3D medical segmentation datasets. To achieve such goal, we have invested significant efforts in standardizing datasets and unifying annotation labels, paving the way for universal segmentation task. For complete list of datasets and download links, we refer the reader to Table 26 and Table 4.
Figure 2: Overview of our dataset for training universal segmentation model, from a wide range of segmentation tasks across various modalities and anatomical regions of the human body, including the brain, head and neck, thorax, spine, abdomen, upper limb, lower limb, and pelvis.
On architecture design and training strategy, we build a universal medical segmentation model, that enables flexible segmentation across a spectrum of medical imaging modalities and anatomies, with text as prompts. Specifically, we adopt knowledge-enhanced representation learning, leveraging textual anatomical knowledge and atlas segmentation of specific anatomical structure to train the text encoder. Through this training process, the visual features of these anatomical structures are aligned with their corresponding text descriptions in the latent space, which is validated to boost the segmentation performance, especially in a long-tail distribution. Later, the text embeddings of anatomy/abnormality are treated as queries in a Transformer-based architecture, iteratively attending the visual features to extract task-specific features for precise segmentation of the queried target. To meet the requirements from different computational resources, we train two models of varying sizes, namely, SAT-Nano and SAT-Pro, and validate the effectiveness of the scaling law in universal segmentation.
On experiment evaluation, we devise comprehensive metrics for universal medical segmentation across various aspects, including region-wise average, organ-wise average, and dataset-wise average. Through extensive internal and external experiments, we demonstrate that:
  • Building on the unprecedented dataset collection, SAT can handle a wide range of downstream segmentation tasks with flexible text prompts, simplifying the conventional solution to train and deploy numerous specialist models. On internal evaluation, SAT-Pro, with only 447M parameters, shows comparable performance to specialist models trained individually on each dataset (72 separate nnU-Nets around 2.2B parameters); While on external evaluation, SAT-Nano, with only 110M parameters, demonstrate even better generalization ability over nnU-Net models.
  • As a foundational segmentation model pre-trained on the large-scale and comprehensive data, SAT-Pro could be further optimized and adapted to specific tasks, balancing the clinical needs between unievrsal solution and specialist models. On most of the datasets, it shows superior performance over the counterpart nnU-Nets.
  • Driven by text prompts, SAT outlines a novel universal segmentation paradigm as opposed to previous efforts in building interactive models, which relies on spatial prompts. This could save tremendous manual efforts from prompting in clinical applications. On performance, we demonstrate it can outperform the state-of-the-art interactive universal segmentation model, namely, MedSAM, on internal and external evaluations, while being robust to targets with ambiguous spatial relationships.
  • The text-prompted feature and universality of makes SAT a powerful out-of-box agents for language model. We show SAT can be seamlessly integrated with any large language models such as GPT-4 [2], automatically providing explainable grounded report in clinical scenarios, or providing grounding annotation for medical data in a scalable way.
2 Results
The objective of our proposed model, termed Segment Anything with Text (SAT), is to develop a universal segmentation framework for 3D medical images, guided by text prompts, to handle various segmentation tasks using a single model. The universality of SAT not only allows easy adaptation to clinical procedures, meeting a wide range of clinical demands, but also makes SAT a versatile tool that bridges linguistic inputs and segmentation outputs. To this end, we have developed a large-scale and comprehensive data collection, termed SAT-DS, covering a broad spectrum of medical image segmentation tasks. It includes 497 anatomical targets across 8 regions and various lesions of the human body, assembled from 72 distinct datasets.
In this section, we will detail the experiment results where SAT is comprehensively evaluated against two representative baselines: nnU-Nets [27], an ensemble of specialist models optimizing configuration and training individually on each dataset; and MedSAM [54], state-of-the-art interactive universal segmentation model driven by box prompts. The evaluation results are present from multiple perspectives, including region-wise, class-wise, and dataset-wise, to gain a deep understanding of our models’ capabilities. Note that class-wise and region-wise evaluations involved averaging the results from different nnU-Nets. For instance, the performance metrics for the ‘brainstem’ in CT images represent the macro average from models trained on datasets like ‘HAN Seg’, ‘PDDCA’, and ‘SegRap2023 Task 1’, all of which include annotations for this anatomical class. Detailed experiment settings can be found in Sec. 4.4.
2.1 Comparison with Specialist Models on Automatic Segmentation
Evaluation on Internal Datasets. In this experiment setting, we aim to evaluate SAT on in-domain samples, i.e.,3D images that follow the same distribution as train set. Specifically, we split each of the 72 datasets in SAT-DS into train and test splits in 8:2 ratio. We train and evaluate two variants SAT-Pro and SAT-Nano in this setting; While for nnU-Nets, we train 72 nnU-Nets specialized on each dataset and report their results as a whole. Table 1 shows the region-wise results on 8 regions of human body, including Brain, Head and Neck, Thorax, Abdomen, Pelvis, Spine, Upper Limb, and Lower Limb, as well as Lesion, in terms of DSC and NSD, respectively. Classes existing in multiple regions are specifically grouped as Whole Body. Notably, SAT-Pro consistently outperforms nnU-Nets in four regions: on Head and Neck, it surpasses nnU-Nets by 5.21 on DSC and 8.27 on NSD; On Thorax, it surpass nnU-Nets by 0.78 on DSC and 2.49 on NSD; On Upper Limb, it surpasses nnU-Nets by 2.8 on DSC and 9.28 on NSD; On Lower Limb, it surpasses nnU-Nets by 1.53 on DSC and 2.48 on NSD. Fig. 3 provides another view on the above results, where it can be seen that SAT-Pro shows slightly lower or comparable segmentation performance to the 72 nnU-Nets, while being of significantly smaller size, approximately 1/5 of the assemble of nnU-Nets. We present more detailed region-wise results in supplementary material Fig. 13 and Fig. 14. We further present class-wise results in supplementary material Sec. 8.3, and dataset-wise results in supplementary material Sec. 8.2.
Table 1: Internal evaluation results of nnU-Nets, SAT-Nano, SAT-Pro and SAT-Pro-FT on different regions of human body and lesions. H&N: head and neck, LL: lower limb, UL: upper limb, WB: whole body, All: average over all the 497 classes. The best results are bolded.
Metric Method Brain H&N UL Thorax Spine Abdomen LL Pelvis WB Lesion All
[…]
SAT-Nano 73.90 74.08 89.58 82.09 75.55 77.12 81.79 79.50 77.80 46.04 75.51
SAT-Pro 77.70 77.29 91.34 86.38 74.73 81.47 84.01 83.01 82.19 51.39 78.81
SAT-Pro-Ft 78.80 78.36 93.78 89.43 82.48 84.11 87.12 86.84 84.24 53.71 81.41
[…]
SAT-Nano 72.95 79.15 90.94 79.68 73.68 67.83 78.42 74.89 75.36 38.17 73.76
SAT-Pro 77.77 82.45 93.56 85.15 72.87 72.94 82.50 78.66 79.61 44.87 77.82
SAT-Pro-Ft 80.43 84.01 95.64 86.38 82.37 78.42 87.25 82.92 82.48 47.74 81.60
Figure 3: Comparison between SAT-Nano, SAT-Pro and nnU-Nets on different regions and lesions. (a) DSC score comparison; (b) NSD score comparison; (c) Model size comparison. SAT-Pro is approximately 5 times smaller than the total size of all the specialist nnU-Nets, while SAT-Nano is merely one twentieth the total size of all the specialist nnU-Nets.
Dataset-specific Adaptation. Built on the unprecedented data collection, SAT can be used as a strong pre-trained model, and adapted towards specific downstream dataset, by further finetuning on the training split of certain dataset. As shown in Table 1, we also report the region-wise results of SAT-Pro-FT. When comparing to the original SAT-Pro, it shows significant improvement on all the regions and lesion classes, especially on Spine (+7.75 on DSC & +9.5 on NSD) and Pelvis (+3.83 on DSC & +4.26 on NSD). Overall, after finetuning on each of the 72 datasets, and compute the average, SAT-Pro-Ft improves 2.6 on DSC and 3.78 on NSD, ultimately outperforms nnU-Nets by 0.87 on DSC, and 3.45 on NSD. The class-wise results are presented in supplementary material Sec. 8.3, and dataset-wise results in supplementary material Sec. 8.2.
Evaluation on External Datasets. In this experiment setting, we aim to evaluate the generalization performance segmentation models on images outside the train set distribution, e.g.,to data from different centers. Note that, this do not involve new classes, as the segmentation targets in human body are relatively limited and fixed. To this end, we use a subset collection of SAT-DS as training data, termed as SAT-DS-Nano, covering 49 datasets, and 429 classes; and take 10 held-out datasets for evaluation. The details of SAT-DS-Nano can be found in Sec. 4.4, and the composition can be found in Table 26. Though we found Although SAT-Pro is generally superior over SAT-Nano in internal evaluation, to save computational cost, we only train and evaluate SAT-Nano in this setting; While to systematically evaluate nnU-Nets, for each held-out dataset, we traverse all the 49 datasets and corresponding models, and evaluate them on all the classes shared between the two datasets. For example, to segment ‘lung’ in COVID-19 CT Seg, we utilize nnU-Nets trained on CT-ORG and LUNA16, as they all involve this class, and then average the results. Overall, 23 out of 49 nnU-Nets are transferred for 80 times and are evaluated as a whole. The transfer details are in supplementary material Fig. 19. To maintain nnU-Nets’ performance, the target dataset are processed consistently with the source dataset in transfer.
Table 2: External evaluation results on each held-out dataset. AA stands for AbdomenAtlasMini1.0 [71] (a very recently released segmentation dataset not included in SAT-DS); CHAOS stands for CHAOS CT; DA stands for DAP Atlas; Liver stands for MSD Liver; Pancreas stands for MSD Pancreas; Prostate stands for MSD Prostate; Spleen stands for MSD Spleen; TS V2 stands for TotalSegmentator V2. Here, SAT-Nano denotes the variant trained with held-out datasets. The best results are bolded. Note that, nnU-Nets and SAT-Nano are fully automatic methods, while MedSAM is interactive. Datasets involved in training MedSAM are excluded and marked /.
Metric Method AA BTCV CHAOS COVID19 DA Liver Pancreas Prostate Spleen TS V2
[…]
SAT-Nano 83.56 76.01 96.69 94.6 61.53 96.58 87.21 72.76 93.68 46.8
Oracle Box 75.26 68.82 72.93 / 54.52 73.13 / / / /
MedSAM (Tight) 88.41 77.94 96.48 / 64.03 95.87 / / / /
MedSAM (Loose) 75.39 56.01 94.15 / 41.07 48.66 / / / /
[…]
SAT-Nano 75.59 71.94 82.35 86.83 57.65 79.46 75.62 44.28 81.71 37.76
Oracle Box 45.92 54.36 28.28 / 44.64 26.31 / / / /
MedSAM (Tight) 78.43 74.4 81.0 / 55.03 77.26 / / / /
MedSAM (Loose) 45.7 36.04 66.87 / 21.47 27.89 / / / /
We report dataset-wise results in Table 2. In general, SAT-Nano outperforms nnU-Nets on 9 out of 10 datasets on DSC score (+2.21), and excels consistently on all datasets on NSD score (+6.73). Notably, SAT-Nano exceeds nnU-Nets significantly on datasets like MSD Prancreas (+13.94 on DSC & +22.99 on NSD), MSD Liver (+1.17 on DSC & +10.82 on NSD) and MSD Prostate (+8.91 on DSC & +1.72 on NSD). We conjecture this is primarily because the dataset-wise customization hinders the generalization of nnU-Nets to external datasets. Note that, based on the universal segmentation capability, SAT-Nano only needs the user to specify the name of the interested region at inference time, while nnU-Nets require cumbersome transfers from models specialized on different datasets and class sets (label space).
2.2 Comparison with Interactive Segmentation Model
In this section, we compare our proposed SAT with MedSAM, As MedSAM is an out-of-the-box universal segmentation methods trained on large-scale data, we directly use it on our data without finetuning. Similar to Sec. 2.1, the comparison with MedSAM is conducted in both internal evaluation and external evalution. However, due to the inconsistent training data, we carefully choose 32 out of 72 datasets for internal evaluation, which are also involved in training MedSAM. Note that, even though these datasets are included in training MedSAM, the train-test splits are not aligned, which means the test set we used might be leaked in training MedSAM. The detailed composition can be found in supplementary material Table 10. And 5 out of 10 held-out datasets are selected for external evaluation, which are excluded in training MedSAM. In both internal and external evaluation, SAT-Pro and SAT-Nano refer to the same models in Sec. 2.1, only the number of evaluation datasets are reduced. We report three results for MedSAM: simulate box prompts based on the ground truth segmentation, with minimum rectangle covering the ground truth (denoted as Tight), i.e., the most accurate prompts; randomly shift each box corner up to 8% of the image resolution (denoted as Loose), i.e., allow errors to some extent; directly use the tight box prompts as prediction (denoted as Oracle Box), i.e., the input baseline of MedSAM.
Evaluation on Internal Datasets. Table 3 shows the region-wise results of all methods. Notably, SAT-Pro consistently outperforms MedSAM on all regions in terms of both DSC ($+7.08$) and NSD ($+10.8$), even if MedSAM is prompted with the most accurate box (Tight). On lesion segmentation, SAT-Pro underperforms MedSAM (Tight), due to the small size of lesion, and the box prompts to MedSAM have already provided very strong prior, as indicated by the results of inputting oracle box, which even shows better performance than MedSAM output on DSC score. We present more detailed region-wise results in supplementary material Fig. 17 and Fig. 18, class-wise results in supplementary material Sec. 8.6 and dataset-wise results in supplementary material Sec. 8.5.
Table 3: Internal evaluation results of SAT-Pro, SAT-Nano and MedSAMs on different regions of human body and lesion. H&N: head and neck, All: average over all the classes. The best results are bolded. Note that SAT-Pro and SAT-Nano are fully automatic methods, while MedSAM is interactive.
Metric Method Brain H&N Thorax Spine Abdomen Limb Pelvis Lesion All
[…]
SAT-Nano 69.05 77.93 79.06 81.88 78.43 77.6 78.31 47.7 76.66
Oracle Box 55.84 85.32 72.85 78.11 71.92 78.96 81.3 67.94 63.57
MedSAM (Tight) 54.35 78.48 73.01 79.09 77.35 80.53 84.42 65.85 75.39
MedSAM (Loose) 14.7 44.68 36.49 45.72 19.79 42.97 51.42 15.71 35.18
[…]
SAT-Nano 75.62 79.44 77.74 82.7 74.96 77.58 72.98 40.38 74.82
Oracle Box 60.55 86.83 66.16 69.86 59.51 70.67 71.59 61.59 48.93
MedSAM (Tight) 68.62 82.37 68.26 73.29 73.39 73.84 74.9 62.28 70.99
MedSAM (Loose) 10.29 43.57 31.61 43.84 16.81 41.54 46.74 11.46 31.50
We further present the qualitative results from two representative examples to illustrate the performance gap between SAT-Pro with MedSAM. In Fig. 4, the upper row shows myocardium segmentations on a CT image. As the myocardium is hard to be referred by the rectangular box prompt, MedSAM generate unsatisfactory prediction which wrongly includes the left heart ventricle surrounded by the myocardium. By comparison, SAT-Pro generate the desirable prediction via the word ‘myocardium‘. It reveals the fundamental dilemma underlying MedSAM: spatial prompts can be ambiguous in complicated situations. The lower row demonstrates colon cancer segmentations on a CT image. The tight box prompt MedSAM receives forms an acceptable segmentation, despite its limitation as a rectangle. While by contrast, the segmentation output from MedSAM is even worse. This confirms our conjecture that MedSAM performs well in lesion targets primarily due to the strong prior provided by the tight box prompt.
Figure 4: Visualization of MedSAM (Tight) and SAT-Pro results on liver (upper row) and colon cancer (lower row) segmentation examples. The ground truth and segmentation masks are painted in red, while box prompts of MedSAM are plotted in blue. The DSC score is calculated in slice-wise manner.
Evaluation on External Datasets. The dataset-wise results are shown in Table 2. Compared with MedSAM (Tight), SAT-Nano achieves higher DSC score on CHAOS CT (+0.21) and MSD Liver (+0.71); and higher NSD score on CHAOS CT (+1.35), DAP Atlas (+2.62) and MSD Liver (+1.8). While for MedSAM (Loose), the performance drops significantly on all datasets, indicating that the model is sensitive to user inputs.
2.3 Ablation Study
As will be illustrated in Sec. 4.2, to build an universal segmentation model driven by text prompts, we devise the visual encoder and decoder to process diverse medical images and perform segmentation; Meanwhile, domain knowledge is injected into the text encoder to provide precise prompts for the target of interest, i.e., the encoding of terminology. In this section, we conduct experiments and discuss the effect of different visual backbones and domain knowledge. To save computational cost, all the experiment in this section are conduct on SAT-DS-Nano dataset.
2.3.1 Effect of Visual Backbone
Figure 5: Evaluations on different visual backbones on SAT-DS-Nano. ‘All’ denote the average scores over all the classes(n=429), including lesion classes. (a) DSC comparison; (b) NSD comparison. U-Net-CPT and U-Mamba-CPT perform very close, while both surpass SwinUNETR-CPT considerably.
Extensive studies have been done discussing the applicability of different visual backbones across segmentation tasks. However, it remains unexplored on which visual backbone suits the need of universal segmentation. In addition to the ConvNet-based U-Net, we consider two alternative backbones, namely, SwinUNETR [20] and U-Mamba [18] to medical segmentation. We configure three SAT-Nano variants with SwinUNETR(107M), U-Mamba(114M) and U-Net(110M) in comparable size: For SwinUNETR, we use the same hyper-parameters as in official implementation [19]; For U-Mamba, we refer to the official implementation [53], reserving configuration of the U-Net, and only add mamba layers at the end of last 3 blocks of the U-Net encoder.
To avoid repeating multimodal knowledge injection on each visual backbones, we use MedCPT [30] for all these variants as text encoder, without loss of generality. MedCPT is a text encoder trained on 255 million in-house user click logs from PubMed [76], and shows state-of-the-art performance on various biomedical language tasks. We expect it to possess domain knowledge and embed terminologies properly as text prompts. Thus, we denote these variants as U-NET-CPT, SwinUNETR-CPT and U-Mamba-CPT. We demonstrate the region-wise evaluation results in Fig. 5. U-Net-CPT outperforms U-Mamba-CPT slightly on both DSC (0.35) and NSD (0.22) scores averaged over all classes. However, both U-Net-CPT and U-Mamba-CPT exceed SwinUNETR-CPT by a significant margin. These observations confirms our choice of U-Net as the visual backbone of SAT. For more detailed comparison on each dataset and class, we refer the reader to supplementary material Sec. 8.9 and Sec. 8.10.
2.3.2 Effect of Text Encoder
Figure 6: Evaluations on different text encoders on SAT-DS-Nano. ‘All’ denote the average scores over all the classes(n=429), including lesion classes. (a) DSC comparison; (b) NSD comparison. Prompted by MedCPT, SAT achieves better performance over BERT-base on average over all the classes. While our pre-trained text encoder outperforms MedCPT consistently across all the regions and lesion.
In this section, we investigate the impact of text encoder for the text-prompted universal segmentation. Specifically, we train three SAT-Nano models with three representative text encoders: the text encoder pre-trained on our multimodal medical knowledge graph, as illustrated in Sec. 4.2.2; the state-of-the-art text encoder for various medical language tasks, e.g., MedCPT; And BERT-Base, a prevalent text encoder in natural language processing but not specifically finetuned on medical corpus. For all the variants, we use U-Net as visual backbone, and denote them as U-Net-Ours, U-Net-CPT, U-Net-BB.
Region-wise Evaluation. As shown in Fig. 6, the performance of U-Net-CPT and U-Net-BB are close, with U-Net-CPT slightly outperfoms U-Net-BB on both DSC (+0.43) and NSD (+0.52) scores averaged over all classes, suggesting that finetuned on medical corpus, MedCPT might hold minor advantage. By contrast, U-Net-Ours surpasses U-Net-CPT consistently on all regions and lesions, leading notable margins on both DSC (+1.54) and NSD (+2.65) scores after average over all classes. This demonstrates the effectiveness of our proposed multimodal knowledge injection. For more detailed comparison on each dataset and class, we refer the reader to supplementary material Sec. 8.7 and Sec. 8.8.
Medical Concept-to-definition Retrieval. To demonstrates advantage our proposed knowledge injection on capturing the critical features of medical terminologies, we also evaluate the three text encoders on concept to definition retrieval task, with the human anatomy knowledge we collect, totaling 6,502 anatomy concepts and definition pairs. We find that the Recall1 (R1) for BERT-Base is merely 0.08%, suggesting it can hardly understand these anatomy concepts or possess almost no domain knowledge; The R1 for MedCPT is 11.19%. Though this is a significant improvement over BERT-Base, MedCPT still struggles to distinguish these concepts; By contrast, our proposed text encoder get 99.18% R1, indicating the knowledge is successfully injected into the text embedding for each anatomy concept.
Knowledge Injection Helps Segmentation of Long-tail Classes. The 429 classes in SAT-DS-Nano follow a typical long-tail distribution, as illustrated in Fig. 7 (a) and (b). The 10 classes with maximum annotations account for 12.75% of all the annotations in SAT-DS-Nano, denoted as ‘head’ classes; On contrast, the 150 classes with minimum annotations only accounts for 3.25%, even though they comprise 34.97% of 429 classes. We compare U-Net-Ours, U-Net-CPT and U-Net-BB on the ‘head’ classes, ‘tail’ classes and the rest ones (denoted as ‘middle’ classes). In Fig. 7 (c), we observe that prompted by our proposed text-encoder, U-Net-Ours achieves the best performance on all three type of classes. On ‘head’ classes, it outperforms the second best variant by 0.71 on DSC and 2.44 on NSD. While on ‘tail’ classes, the improvement is even more significant: 1.4 on DSC and 3.23 on NSD. Meanwhile, U-Net-CPT is better than U-Net-BB on ‘tail’ but slightly underperform on ‘head’. These results suggest that our proposed knowledge injection can effectively boost the segmentation performance on ‘tail’ classes.
Figure 7: The impact of domain knowledge on a long-tail distribution. (a) The annotation number of all the 429 classes in SAT-DS-Nano, sorted from highest to lowest. (b) The proportion of ‘head’, ‘middle’ and ‘tail’ in class number and annotation number. The DSC and NSD comparison on ‘head’, ‘middle’ and ‘tail’ classes.
2.4 Qualitative Results – SAT as an Interface Between Language and Segmentation
Benefiting from the universality and text-driven features of SAT, it can be flexibly applied as an interface between natural language and segmentation. Specifically, we utilize GPT-4 [2] to directly extract the anatomical targets of interests from real clinical report and prompt SAT to segment them on the clinical image. In Fig. 8 and Fig. 9, we demonstrate the zero-shot performance of SAT-Pro on four cases randomly selected from clinical practice: abdominal MR examination, chest CT examination, abdominal CT examination, and lumbar spine CT examination. The basic information of patients, the report written by doctors, and targets of interests detected by GPT-4 are attached to each cases. As can be seen, the targets in reports can be well detected by language model (GPT-4) and commendably segmented by SAT-Pro: In Fig. 8, Liver, Pancreas, Kidneys, Gallbladder and Spleen are segmented in both abdominal CT and MR images; While in Fig. 9, segments of right lung, heart and bronchi are segmented in thoracic CT image, lumbar vertebrae L1 to L5 are segmented in spinal CT image.
Figure 8: Zero-shot transfer to real clinical data. Top: A case from abdominal CT examination, no abnormal is observed from the CT image; Bottom: A case from abdominal MR examination, no abnormal is observed from the MR image; Basic information of patients, predefined targets of interests, and final reports written based on the segmentation results were appended to each cases.
Figure 9: Zero-shot transfer to real clinical data. Top: A case from chest CT examination, the patient is taking a follow-up examination after treatment for synovial sarcoma with pulmonary metastasis; Bottom: A case from lumbar vertebrae CT examination, some abnormalities are observed on the image; Basic information of patients, predefined targets of interests, and final reports written based on the segmentation results were appended to each cases.
3 Discussion
Developing specialist models on individual tasks has been the dominant solution for medical image segmentation for years [27, 15, 44, 95, 102, 79, 99, 21, 11, 19, 93, 98, 12, 101]. Yet the heterogeneity of medical images has long called for effective and flexible universal medical segmentation model to meet clinical needs. To this end, we proposed a universal segmentation model, trained on unprecedented dataset collection, and driven by knowledge enhanced text prompt. In this section, we discuss the significance, potential and limitations of the proposed SAT, based on our experiment results in Sec. 2.
A Universal Model is Worth 72 Specialist Models. Through comprehensive internal evaluation from several aspects, SAT-Pro has demonstrated comparable results to an assemble of 72 specialized nnU-Nets, and even surpass them on lots of regions and classes. While on external evaluation, SAT-Nano exhibits superior generalization ability than nnU-Nets. Therefore, despite the diversity of image and segmentation targets from different clinic scenes, SAT can flexibly handle them with just one model. SAT along with the unprecedented dataset collection and standardisation procedure, represent an important step forward for universal segmentation model in medical scenario.
SAT as a Strong Pre-trained Segmentation Model. When fine-tuned on specific dataset, SAT-Pro-Ft can further improve the performance, achieving superior results over the counterpart nnU-Nets. Thus, it naturally acts a foundational segmentation model, pre-trained on large-scale and diverse data, holding potentials to be adapted to specific tasks, balancing the clinical requirements between universal solution and specialist model.
Text As Prompt for Segmentation. As an automatic method prompted by text, SAT outlines a new path apart from recent works in interactive universal segmentation [54, 17, 87, 87]. In both internal and external evaluation, SAT saves time from laborious spatial prompting, showing superior scalability in deployment, provides semantic and robust results regardless of manual errors in drawing box prompts.
SAT Can Be Used As An Agent to Bridge Language and Segmentation. In Sec. 2.4, we show that SAT can be applied to segment targets extracted by language model from clinical report, providing explainable grounded report in clinical scenarios. This demonstrate the potential of SAT as a powerful out-of-the-box agent, seamlessly integrated into any large language models via text prompts. As grounding is an indispensable procedure for building generalist medical artificial intelligence [63], we expect the potential and significance of SAT to extend beyond a simple clinical tools, for example, providing grounding annotations for medical datasets in a scalable way [97].
Scaling Models Lead to Better Performance. Increasing model capacity has been demonstrated to effectively enhance model’s performance in large language models [92, 84], yet such scaling law remains to be verified in segmentation tasks. In this work, we build SAT-Nano (110M) and SAT-Pro (447M). In both region-wise and class-wise evaluations, SAT-Pro shows clear performance boost over SAT-Nano, outperforming the latter on most region and classes. This implies that scaling law is also applicable to universal medical segmentation, indicating a promising way to continuously improve the performance.
Domain Knowledge Can Boost Segmentation Performance. The impact of domain knowledge, especially multi-modal knowledge on human anatomy remains under-explored for segmentation task. In this work, we propose a multi-modal knowledge graph on human anatomy, and via knowledge injection, build a text encoder prompting the segmentation task. In Sec. 2.3.2, we demonstrate how it can enhance the segmentation performance, especially on these ‘tail’ classes. As we expect the segmentation targets in human body to continuously grow in the future, it will raise higher demand on generating informative and distinguishable text prompts.
Limitations. As an exploratory work in this field, though demonstrate impressive and promising results, we find certain limitation of SAT:
  • The performance of SAT-Pro still lags behind nnU-Nets in some region including Brain, Spine and Abdomen, and many classes. In future work, we will continue scaling up model size, to further close the gap and even outperforms the counterparts specialist models.
  • SAT currently only supports text as prompts, thus not intended for scenes with human interaction requirements. For example, manual refinement on the segmentation results. We plan to incorporate more interaction interfaces, for example, boxes, scribbles.
  • The distribution of SAT-DS is unbalanced, leading to the limitation on capability of SAT. For instance, SAT-DS contains 189 classes for MR images, while 137 classes are on ‘Brain’ region. Comprehensive annotation for MR images still lags behind CT images in the scope of public datasets.
  • Due to the imbalance between different segmentation datasets, the long-tail distribution in assembled dataset collection remain challenging for building an universal segmentation method. As demonstrated in Fig. 7 (c), the performance gap between ‘head’, ‘middle’ and ‘tail’ classes is still significant.
4 Method
4.1 Dataset Construction
Toward building a universal segmentation model for medical imaging, with text as prompts, we collect two types of data in this work: medical domain knowledge to train the text encoder, and medical segmentation data, which includes both images and segmentation masks. In this section, we will describe them sequentially.
4.1.1 Domain Knowledge
To acquire textual knowledge, we mainly exploit two sources of domain knowledge: the Unified Medical Language System (UMLS) [^1] https://www.nlm.nih.gov/research/umls, a comprehensive medical knowledge graph consisting of concept definitions and their interrelations; search engines, which are prompted to organize knowledge into a graph of the same format, specifically refined for the human anatomy corpus. Regarding visual knowledge, we have compiled 72 medical segmentation datasets, creating an atlas that covers over 497 anatomical structures of the human body. All sources of knowledge are systematically categorized, encompassing anatomies and concepts from all regions of the human body. Two examples of these sources are illustrated in Fig. 10 (a) and (b).
UMLS. Unified Medical Language System (UMLS) [10] is a knowledge source of biomedical vocabulary developed by the US National Library of Medicine [^2] https://www.nlm.nih.gov. It integrates a wide range of concepts from more than 60 families of biomedical vocabularies, each equipped with a Concept Unique Identifier (CUI) and definition. It also contains the relations among these concepts. Following [96], we extract 229,435 biomedical terminologies and definitions, as well as 1,048,575 relationship triplets, composing a knowledge graph on these terminologies.
Domain Knowledge from Search Engine. Although UMLS is widely acknowledged and adopted as a general medical knowledge corpus [91, 96, 42, 100], it lacks a fine-grained description of anatomy concepts critical for segmentation tasks. For example, for ‘liver’, the definition is ‘A large lobed glandular organ in the abdomen of vertebrates that is responsible for detoxification, metabolism, synthesis, and storage of various substances.’, which erases the morphological features and focuses on functionality. Meanwhile, more comprehensive knowledge on human anatomy may be scattered across various authoritative websites online, e.g., Wikipedia [^3] https://www.wikipedia.org, ScienceDirect [^4] https://www.sciencedirect.com/browse/journals-and-books?subject=medicine-and-dentistry, etc.. To harvest such knowledge, we select 6,502 anatomy concepts, and prompt an search engine  [^5] https://you.com to retrieve and summarize definitions for them. We use the following prompt template: {mdframed}[backgroundcolor=gray!10, linewidth=0pt] Definition of xxx. Include the location, shape, appearance, structure and spatial relations to other anatomy structures. No need to include functionality. End with ‘END’. For illustration, the search engine referred to authority websites including Columbiasurgery [^6] https://columbiasurgery.org/liver/liver-and-its-functions, Hopkins Medicine [^7] https://www.hopkinsmedicine.org/health/conditions-and-diseases/liver-anatomy-and-functions and summerized definition for ‘liver’ as: ‘A large organ found in the upper right quadrant of the abdomen, it stands as the largest gland within the human body, with a weight of about 1.5 kilograms. This structure exhibits a reddish-brown hue and is cone or wedge-shaped … …’. While constructing the knowledge graph, we also adopt GPT4 [2] to extract 38,344 relations between anatomical structures in the generated information-dense definitions with the following prompt: {mdframed}[backgroundcolor=gray!10, linewidth=0pt] This is the description of xxx. Please help me find its relations with other anatomical structures in radiological images. Summarize them with the template: Relation: xxx (relational preposition), Anatomical structure: xxx (name of another anatomical structure).
For example, “Relation: situated below, Anatomical structure: xxx”, “Relation: connected to (via xxx), Anatomical structure: xxx” … …
Visual Atlas. Segmentation datasets naturally provide visual features for anatomy concepts corresponding or complementary to the textual description, such as the texture, spatial location, shape and so on. Details on our collected segmentation datasets are described in Sec. 4.1.2. Here, we use them as a large-scale and diverse visual atlas library, and link the visual regions to corresponding concepts in textual knowledge graph, bridging the knowledge between visual and language modality.
In summary, by mixing these data, we construct a multimodal medical knowledge tree. As demonstrated in Fig. 10 (c), the concepts (including both anatomical structures and lesions) are linked via the relations and further extended with their definitions, containing their characteristics. Additionally, some are further mapped to the visual atlas, demonstrating their visual features that may hardly be described purely by text.
Figure 10: The medical knowledge used in the visual-language pretraining. (a) Segmentation datasets provide an atlas for extensive anatomy concepts. In this example, atlas segmentation is marked with different colors. (b) Knowledge generated from UMLS and search engines encompasses a broad array of concept-definition pairs and extensive relationships. (c) By integrating all collected knowledge sources, a medical knowledge tree is constructed. All definitions are partially displayed for conciseness. Definition and relation denoted with * are derived from the search engine, otherwise from UMLS.
4.1.2 Segmentation Dataset
To train our segmentation model with the ability to handle segmentation tasks of different targets, across various modalities and anatomical regions, we collect and integrate 72 diverse publicly available medical segmentation datasets, totaling 22,186 scans including both CT and MRI and 302,033 segmentation annotations spanning 8 different regions of the human body: Brain, Head and Neck, Upper Limb, Thorax, Abdomen, Plevis, and Lower Limb. The dataset is termed as SAT-DS. More details are present in Table 26 and Table 4. Note that, some public datasets are not mutually exclusive, e.g., KITS23 and KITS21 [24], we thus only collect the latest version, to avoid redundancy and potential data leakage in train-test split.
Before mixing these datasets for training, two challenges remain: (i) the anatomical targets from each dataset must be integrated into a unified annotation system. The clinic demands beneath each dataset collection might be different, resulting in different annotation standards and granularity. Meanwhile, since most datasets are annotated for training specialist models like nnU-Net [27], precise and consistent terminology or expression for anatomical targets is often ignored. Therefore, a unified label system is demanded to avoid potential contradictions when training on mixed datasets. (ii) some critical image statistics, such as intensity distribution and voxel spacing vary from dataset to dataset, hindering the model from learning consistent image representations across datasets. In the following, we present details for dataset integration and pre-processing, and how we address the abovementioned challenges.
Figure 11: Statistics of  SAT-DS across different anatomical regions. (a) Annotation num of some representative classes in each anatomical region; (b) Number of classes in each anatomical region; (c) Number of annotations in each anatomical region. LC/HC: laryngeal/hypopharyngeal cancer, WHM: white matter hyperintensities, PV&SV: portal vein and splenic vein, IVC: inferior vena cava, Thoracic V: thoracic vertebrae, Cervical V: cervical vertebrae, Lumbar V: lumbar vertebrae.
[…]
Dataset Pre-processing. As properties of each dataset may greatly impact the training of the segmentation network [27], such as intensity distribution and voxel spacing, we deliberately apply some normalization procedures to all the datasets to ensure uniformity and compatibility between them. Firstly, all the images are reoriented to specific axcodes, respaced to a voxel size of $1\times 1\times 3\leavevmode\nobreak\ mm^{2}$ and cropped to non-zero region. Secondly, we apply different intensity normalization strategies to CT images and MRI images, specifically, for CT images, intensity values are truncated to $[-500,1000]$ and applied z-score normalization. For MRI images, intensity values are clipped by $0.5\%$ and $99.5\%$ of the image, and then z-score normalized. During training, we randomly crop the image patch with a fixed size of $288\times 288\times 96$. Random zoom-in, zoom-out, and intensity scaling are applied for data augmentation.
Dataset Statistics. After integrating datasets, we derive a segmentation data collection that covers 497 segmentation classes, far outpacing each single dataset in both diversity and scale. We divide human body into eight regions, and classify each class into them manually. Fig. 11 (b) and (c) show the distribution of classes and annotations across different human body regions. We further show the distribution of some example classes in each region in Fig. 11 (a). The extensive range of categories and regions lays the foundation for the SAT’s wide application scenarios. The data collection is more than fourth times the size of the largest dataset (BraTS2023-GLI) in terms of volume number, and nearly triple the most comprehensive dataset (DAP Atlas) in terms of the class number.
Discussion. In the process of building SAT-DS, we merge a wide range of segmentation tasks, and establishing a unified label system by using natural language/text. Generally speaking, there are three advantages to do this: (i) natural language is powerful and discriminative, that enables better differentiation of the medical terminologies in the language embedding space; (ii) as shown in previous work [91, 96, 42, 100], knowledge-enhanced representation learning for the text encoder demonstrates promising performance, allowing to learn the implicit or explicit relationships between these segmentation targets. For example, segmenting a specific lobe of the liver requires the exact segmentation of liver as an organ in abdomen cavity, and shall be facilitated by referring to other parts of the liver. Therefore, establishing such connections via systematic medical knowledge shall be beneficial. (iii) text prompt can be given automatically without any human intervention, for instance, from large language model. This would pave the way for building a universal segmentation model that can be flexibly integrated into foundation models for generalist medical artificial intelligence, as a powerful grounding tool.
4.2 Universal Segmentation Prompted by Text
Towards building up our universal segmentation model by text prompt, we consider two main stages, namely, multimodal knowledge injection and universal segmentation training. In the following, we start by defining the problem scenario (Sec. 4.2.1). Then, we show how to structure multimodal medical knowledge and inject it into a text encoder (Sec. 4.2.2). Lastly, we employ the text encoder to guide our universal segmentation model training on SAT-DS dataset (Sec. 4.2.3).
4.2.1 Problem Formulation
Assuming we have a segmentation dataset collection, i.e., $\mathcal{D}=\{(x_{1},y_{1};T_{1}),...,(x_{K},y_{K};T_{K})\}$, where $x_{i}\in\mathbb{R}^{H\times W\times D\times C}$ denotes the image scan, $y_{i}\in\mathbb{R}^{H\times W\times D\times M}$ is the binary segmentation annotations of the anatomical targets in the image, and $T_{i}=\{t_{1},t_{2},...,t_{M}\}$ denote the corresponding medical terminology set. The universal segmentation task can be formulated as:
$\hat{y}_{i}=\Phi_{\text{SAT}}(\Phi_{\text{visual}}(x_{i}),\Phi_{\text{text}}(T_{i})),$ (1)
where $\Phi_{\text{visual}}$ is a visual encoder, $\Phi_{\text{text}}$ is a text encoder, $\Phi_{\text{SAT}}$ is a universal segmentation model and ideally, $x_{i}$ can be an image scan from any modality of the anatomical region, $T_{i}$ can contain any text-wise medical terminology of interest.
4.2.2 Multimodal Knowledge Injection
Here, we aim to inject rich multimodal medical knowledge into the visual and text encoders. The section starts from the procedure for structuring the multimodal medical knowledge data, and further presents details to use them for visual-language pre-training.
Multimodal Domain Knowledge Construction. As shown in Fig. 12 (a), the data from UMLS, search engine, and segmentation datasets can be aggregated into two formats:
  • Textual Medical Concept Pair. For text-only knowledge, each concept $t_{i}$ is associated with a definition $p_{i}$, constructing pairs of text $(t_{i};p_{i})$. We also derive a knowledge graph that connects the medical concepts through abundant triplet relationships $(t_{i},r_{ij},t_{j})$. This graph can be alternatively seen as a specialized text pair, $(t_{i}+r_{ij};t_{j})$ or $(t_{i};r_{ij}+t_{j})$, where ‘$+$’ refers to string concatenation. In this way, we can thus unify the two kinds of textual knowledge.
  • Visual Medical Concept Pair. To align with the segmentation task, we gather pairs consisting of a concept (can be either an anatomical structure or lesion) and its image atlas. Note that, multiple pairs could be extracted from a single image. These pairs share a similar format to the segmentation data, denoted as $(x_{i},y_{i};t_{i})$, where $x_{i}$ and $t_{i}$ are consistent with their definition in Sec. 4.2.1 and $y_{i}\in\mathbb{R}^{H\times W\times D\times 1}$ is a binary segmentation mask for $t_{i}$.
In summary, all the knowledge can either be represented as pure text description, e.g., $t_{i}$, $p_{i}$, $t_{i}+r_{ij}$, $r_{ij}+t_{j}$, or atlas segmentation $(x_{i},y_{i})$, and paired further.
Figure 12: Overview of SAT. (a) We inject multi-modal medical knowledge into knowledge encoders via contrastive learning. The knowledge is in different formats: atlas segmentation, concepts(terminologies), definitions, and relationships between concepts. We devise visual and text encoder to embed them; (b) We train a universal segmentation network based on the text prompts from the pre-trained text encoder. It is capable of segmenting a wide range of targets for image scans from different modalities and anatomical regions.
Visual-Language Pre-training. As shown in Fig. 12 (a), for pure text description, we encode them with a BERT [33] pre-trained on PubMed abstract [40]:
$z=\Phi_{\text{text}}(\textbf{t}),\text{ }\textbf{t}\in[t_{i},p_{i},t_{i}+r_{ij},r_{ij}+t_{j}],\text{ }z\in\mathbb{R}^{d},$ (2)
where $d$ refers to the feature dimension. For visual concepts, we adopt the visual encoder $\Phi_{\text{visual}}$. Given the excellent robustness and performance of U-Net [27, 77], we apply a standard 3D U-Net encoder to extract multi-scale image embeddings:
$V_{i}=\{v_{i1},v_{i2},...,v_{iS}\}=\Phi_{\text{visual}}(x_{i}),\text{ }v_{is}\in\mathbb{R}^{H_{s}\times W_{s}\times D_{s}\times d},$ (3)
where $V_{i}$ is the multi-scale feature maps from U-Net encoder layers, and $H_{s},W_{s},D_{s},d_{s}$ are the spatial resolutions and channel width at different layers. We further average ROI pooling on these feature maps respectively, based on the down-sampled segmentation mask fitting resolutions at different layers. The concatenated pooled features can thus be treated as representation of the anatomical target on this image, containing multi-scale visual clues for it.
$z=\mathcal{F}_{\text{pooling}}(\Phi_{\text{visual}}(x_{i});y_{i}),\text{ }z\in\mathbb{R}^{d}.$ (4)
We train the visual encoder and text encoder by maximizing the similarities between all positive knowledge pairs, linked by text prompts (medical terminology names), as shown in Fig. 12. Specifically, given $(x_{i},y_{i};t_{i}),(t_{i};p_{i}),(t_{i}+r_{ij};t_{j}),(t_{i};r_{ij}+t_{j})$, for simplicity, we denote all the encoded features as $z$, regardless of their modality format. For a batch of $N$ pairs $\{(z_{1},z_{1}^{\prime}),...(z_{N},z_{N}^{\prime})\}$, we have:
$\mathcal{L}_{\text{knowledge}}=-\frac{1}{N}\sum^{N}_{i=1}(\log\frac{\exp(z_{i}\cdot z_{i}^{\prime}/\tau)}{\sum^{N}_{k=1}\mathbbm{1}_{i\neq k}\exp(z_{i}\cdot z_{k}^{\prime}/\tau)}+\log\frac{\exp(z_{i}\cdot z_{i}^{\prime}/\tau)}{\sum^{N}_{k=1}\mathbbm{1}_{i\neq k}\exp(z_{k}\cdot z_{i}^{\prime}/\tau)}),$ (5)
which is a typical contrastive learning pipeline, with $\tau=0.07$ as temperature.
Discussion. In general, by maximizing the similarities between positive textual and visual feature pairs, we force the text encoders to construct neural representations for medical concepts from two aspects: (i) through the well-established knowledge graph in text form, the text encoder enables to encode relationships between concepts in the latent space; (ii) the model captures the characteristics of anatomical structures and lesions via both visual atlas segmentations and detailed definitions. Therefore, in contrast to the one-hot labeling that treats each anatomical target as of being independent, such continuous neural representation shall provide more helpful guidance for the segmentation task.
4.2.3 Universal Segmentation Training
With the pre-trained visual and text encoder, we now continue the procedure for building the universal segmentation model with text as prompts. Fig. 12 (b) demonstrates the overall framework. Specifically, apart from the pre-trained visual and text encoders, the segmentation model consists of three more components: a visual decoder $\Phi_{\text{dec}}$, a query decoder $\Phi_{\text{query}}$, and a mask generator.
Although a sample in the segmentation dataset collection $(x_{i},y_{i};T_{i})$ may contain multiple annotations, i.e., $T_{i}=\{t_{1},t_{2},...,t_{M}\}$, for simplicity, we describe the segmentation procedure for one structure ($t_{i}$).
Terminology Encoding. We use the pre-trained text encoder to generate neural embedding for one anatomical target as a text prompt for segmentation:
$z_{i}=\Phi_{\text{text}}(t_{i}),\text{ }z_{i}\in\mathbb{R}^{d}.$ (6)
Note that, after pre-training the text encoder with domain knowledge injection, $z_{i}$ should contain both the textual background information and visual information from atlas samples.
Visual Encoding. We adopt the pre-trained visual encoder as explained in Sec. 4.2.2, Equa. 3, and continue training it.
Visual Decoding. In the visual decoder, the feature maps from the encoder are gradually upsampled with skip connections, effectively following the U-Net architecture [27, 77], ending up with per-pixel dense feature:
$u_{i}=\Phi_{\text{dec}}(V_{i}),\text{ }u_{i}\in\mathbb{R}^{H\times W\times D\times d^{\prime}},$ (7)
where $d^{\prime}$ is the dimension for the per-pixel dense feature after recovering to the original resolution.
Query Decoder. Although a general representation of the anatomical target is derived from the pre-trained text encoder with a text prompt, visual variations may still exist from patient to patient, we thus insert a transformer-based query module to further enhance the text prompts with visual clues. In practice, this module consists of 6 standard transformer decoders [86], that treats text embedding as query, and the pooled multi-scale visual features from the U-Net encoder as key, values, formulated as:
$q_{i}=\Phi_{\text{query}}(V_{i},z_{i}),\text{ }q_{i}\in\mathbb{R}^{d}.$ (8)
Where $z_{i}$ is consistent with $z$ in Equa. 4. Therefore $q_{i}$ can be seen as an adapted representation of the anatomical target in specific image scan $x_{i}$.
Mask Generator. By conducting pixel-wise dot-product between the representation of the anatomical target and the fine-grained per-pixel embedding, we can acquire a per-pixel segmentation:
$\hat{y}_{i}=g(q_{i})\cdot u_{i},\text{ }\hat{y}_{i}\in\mathbb{R}^{H\times W\times D},$ (9)
where $g(\cdot)$ is a feed-forward layer projecting $q_{i}$ to a consistent dimension with the dense feature map $u_{i}$.
Training Objective. Following [27], we adopt a loss function as the sum of binary cross-entropy loss and dice loss. For a sample with $M$ classes and $C$ voxels, we denote $p_{\text{c},\text{m}}$ and $s_{\text{c,m}}$ the prediction and ground-truth for $c$-th pixel respectively on class $m$, the loss is:
$\mathcal{L}=\underbrace{-\frac{1}{M}\sum_{\text{m=1}}^{M}\frac{1}{C}\sum_{\text{c=1}}^{C}p_{\text{c,m}}\cdot\log s_{\text{c,m}}}_{\text{Binary Cross Entropy Loss}}+(\underbrace{1-\frac{2\sum_{\text{i=1}}^{M}\sum_{\text{c=1}}^{C}p_{\text{c,m}}\cdot s_{\text{c,m}}}{\sum_{\text{m=1}}^{M}\sum_{\text{c=1}}^{C}p_{\text{c,m}}^{2}+\sum_{\text{m=1}}^{M}\sum_{\text{c=1}}^{C}s_{\text{c,m}}^{2}}}_{\text{Dice Loss}})$ (10)
4.3 Implement Details
We encounter several challenges in training on the combined large number of heterogeneous medical datasets in 3D format. In this section, we provide details on the adopted training strategies.
Progressive Multimodal Knowledge Injection. For the stability of training, we implement the multimodal knowledge injection procedure introduced in Sec. 4.2.2 progressively. At the beginning, the text encoder is pretrained on the text knowledge via contrastive learning, i.e., the textual medical concept pairs. The maximal sequence length after tokenization is 256, for even longer text input, we apply random truncation to fully exploit the knowledge in the long text. After convergence, to align text representations and visual features, we apply contrastive learning between the finetuned text encoder and the visual encoder on the visual medical concept pair.
Pre-processing Segmentation Dataset. We take voxel spacing as $1\times 1\times 3\ mm^{2}$ and patch size $288\times 288\times 96$, based on two empirical considerations: (i) when mixing the various datasets, scans with a wide range of voxel spacings ought to be normalized before processing in the same convolutional network. While resampling to larger voxel spacing may lose information, smaller voxel spacing will generate artifacts; (ii) a larger receptive field generally ensures better segmentation performance for most targets, however, increasing patch size will result in higher computational cost. We strike a balance between them based on the computational resources in use.
Balancing Segmentation Datasets. To balance between all datasets, we set the sampling strategy for each scan based on two intuitions respectively: (i) training case number varies significantly from dataset to dataset, which should be alleviated. We follow  [85] and set the sampling weight of all scans in a dataset as the inverse proportion to $\sqrt{N}$, $N$ is the number of training cases in the dataset; (ii) scans with larger annotation areas or more annotated classes should be sampled more as they are often harder to learn. Thus, we repeat such scans for $R$ times in the sampling pool, where $R=\frac{S_{roi}}{288\times 288\times 96}\times\frac{M}{32}$. $S_{roi}$ is size of the annotated area in a image scan, namely, its foreground area; $288\times 288\times 96$ is the patch size to crop; $c$ is the number of annotated classes on it; and 32 is the maximal number of text prompts in a batch.
Balancing Segmentation Classes. While cropping the image scan, it’s a common practice [27] to over-sample foreground crops, i.e., crops containing at least one segmentation target. However, weighting these crops evenly may ignore the unbalanced spatial distribution of segmentation targets. For example, in large scans with numerous annotations, tiny targets are harder to sample and thus may be ignored by the model. Thus, in foreground oversampling, we give more weight on regions with more segmentation targets.
Difference between SAT-Pro and SAT-Nano. We devise two variants of SAT with different model sizes in visual backbone. For SAT-Nano, we adopt a U-Net with 6 blocks in-depth, each block has 2 convolutional layers and $3\times 3$ size each kernel. The channel widths for each stage are $[64,64,128,256,512,768]$; SAT-Pro share the same architecture with SAT-Nano, except that each block consists of 3 convolutional layers, and the channel widths are doubled to $[128,128,256,512,1024,1536]$.
Other Hyperparameter Details. The query decoder is a 6-layer standard transformer decoder with 8 heads in each attention module. Feature dimensions of a text prompt $d=768$ and per-pixel embedding $d^{\prime}=64$. To unify the features from visual backbone variants with varying channel widths, they are projected to $d=768$ with different feed-forward layers, and input as key and value to query decoder. For images with multiple segmentation targets, we set the maximal text prompts sampled in a batch of up to 32. A combination of cross-entropy loss and dice loss is applied as supervision at training time. We use AdamW [49] as the optimizer with cosine annealing schedule, maximal $lr=1\times 10^{-4}$, 10000 steps for warm-up. The SAT-Nano is trained on 8 NVIDIA A100 GPUs with 80GB memory, using maximal batch size 2; while the SAT-Pro is trained on 16 NVIDIA A100 GPUs with 80GB memory, using maximal batch size 1.
4.4 Evaluation Settings
4.4.1 Evaluation Datasets
To strike a balance between extensive experiments and computational cost, we utilize two collection of datasets:
  • SAT-DS. As describe in Sec. 4.1.2, this contains all the 72 datasets, 497 classes from all human body regions, 22,186 image scans and 302,033 segmentation annotations.
  • SAT-DS-Nano. A subset of SAT-DS, including only 49 datasets, 13,303 images and 151,461 annotations. Note that SAT-DS-Nano also covers 429 classes from all human body regions, adequate to evaluate universal segmentation task.
The detailed composition of SAT-DS and SAT-DS-Nano can be found in supplementary material Table 26 and Table 27. As there is no existing benchmark for evaluating the universal segmentation model, we randomly split each dataset into 80% for training and 20% for testing: (i) datasets may share the same images but with different classes. For example, Couinaud Liver provides fine-grained liver segmentation on a subset of MSD Hepatic Vessel. We carefully split the Couinaud Liver to make sure the test set will not be leaked in the train set of MSD Hepatic Vessel; (ii) scans of the same patient but different modalities are treated as different samples during training and evaluation. For example, MSD Prostate contains T2 and ADC scans of each patient. However, they share the same structure on the image. To avoid potential data leaking, we carefully split such datasets by patient id. Note that when involving segmentation datasets in the visual-language pretraining, we only use the training data to avoid potential data leaking. For datasets involved in SAT-DS-Nano, we keep their splits the same as in SAT-DS. The download link for each datasets can be found in Sec. 7, and we have released our dataset processing code and train-test splits to the research community for reproduction and benchmarking.
4.4.2 Baseline
Specialist Model. We take nnU-Net [27] as a strong baseline for comparison, which specializes in a series of configurations for each dataset. For a comprehensive evaluation, we train one nnU-Net model on each of the datasets. Note that, following [90], we split Totalsegmentator into 6 subsets and train 6 different nnU-Net models. Similarly, datasets such as CHAOS with both MRI and CT images are treated as two different datasets. When training nnU-Nets on each dataset, we adopt a multi-class segmentation setting and deliver the masks of all categories in this dataset at once. We derive the optimal network architecture and pre-processing pipeline with the default setting of nnU-Net. The detailed network design is shown in supplementary material Table 20 and Table 21. In summary, we train an assemble of 72 nnU-Nets, that are customized on each dataset and serve as the competitive baseline for evaluation. We adopt the latest official implementation of nnU-Net v2888https://github.com/MIC-DKFZ/nnUNet/ in practice. In zero-shot evaluation,
Interactive Segmentation Model. We take MedSAM [54] as a representative interactive universal segmentation model as baseline. MedSAM finetunes SAM [34] on 86 segmentation datasets, and supports 2D medical image segmentation with box prompts. We follow the official implement [^9] https://github.com/bowang-lab/MedSAM to process and infer image slice by slice, and calculate the metrics on the finally stacked 3D prediction. For each single target on a slice, to simulate box prompts towards it, we both take the minimum rectangle containing the ground truth segmentation (Tight), and following the official data augmentation procedure, randomly shift each corner up to 8% of the whole image resolution (Loose). In addition, we consider directly use the tight box prompts as prediction (Oracle Box).
4.4.3 Evaluation Protocols
Given our goal is to develop a universal medical segmentation model, this provides opportunities to evaluate novel perspectives in addition to the traditional evaluation per dataset. Specifically, we conduct the internal evaluations from three dimensions:
  • Class-wise Evaluation. As SAT is capable of segmenting a wide range of anatomical targets across the human body, we merge the results from the same classes across datasets to indicate the performance on each anatomical target. Specifically, we follow macro-average method: for a class annotated in multiple datasets, we first calculate its average scores within each dataset, and then average them over all datasets. Note that, the same anatomical structures or lesions from different modalities are treated as different classes in this work, e.g., liver in both CT and MRI images.
  • Region-wise Evaluation. In general, anatomical structures from the same human body region are closely connected and more likely to be involved in diagnosis within the same hospital department. Here, we consider the region-wise evaluation: based on class-wise evaluation, we merge results from all classes in the same body region, as to indicate the general performance in this region. For classes existing in multiple regions, we classify them into ‘Whole Body’ category. In addition, we report results for lesions classes independently as a category ‘lesion’, instead of merging them into specific regions.
  • Dataset-wise Evaluation. Results of the classes within the same dataset are averaged to indicate the performance on this dataset. This is the same as the conventional evaluation protocol of specialist segmentation models trained on a single dataset.
4.4.4 Evaluation Metrics
We quantitatively evaluate the segmentation performance from the perspective of region and boundary metrics [58], e.g., Dice Similarity Coefficient (DSC) and Normalized Surface Distance (NSD) respectively.
DSC. Dice Similarity Coefficient (DSC) is a standard region-based metric for medical image segmentation evaluation. It measures the overlap between model’s prediction $P$ and ground truth $G$, formally defined as:
$DSC(P,G)=\frac{2|P\bigcap G|}{|P|+|G|}.$ (11)
NSD. Normalized Surface Distance (NSD) [64] is a boundary-based metric that measures the consistency at the boundary area of the model’s prediction $P$ and ground truth $G$, which is defined as:
$NSD(P,G)=\frac{|\partial P\bigcap B_{\partial G}|+|\partial G\bigcap B_{\partial P}|}{|\partial P|+|\partial G|},$ (12)
where $B_{\partial P}=\{x\in\mathbf{R}^{3}|\exists\hat{x}\in\partial P,||x-\hat{x}||\leq\tau\}$ and $B_{\partial G}=\{x\in\mathbf{R}^{3}|\exists\hat{x}\in\partial G,||x-\hat{x}||\leq\tau\}$ are the boundary area of model’s prediction and ground truth at a tolerance $\tau$, respectively. We set $\tau$ as 1 in the experiments.
5 Conclusion
In this paper, we promote the progress of medical universal segmentation with text as prompt and knowledge enhancement. We build up the largest and most comprehensive 3D medical segmentation dataset, and the first multi-modal knowledge tree for human anatomy. In contrast to existing segmentation methods that requires training and deploying a series of specialized models, or limited to box/point prompts, our model, SAT, enables text-driven universal segmentation on 3 radiological modalities, 497 anatomical structures or lesions, across 8 main human body parts in just one model, demonstrates great efficiency and flexibility in application. We proposed SAT-Nano (110M parameters) and SAT-Pro (447M parameters), demonstrating comparable performance to 72 specialist nnU-Nets, with around 2.2B parameters trained individually on each datasets, and showing even superior generalization ability on external datasets. Compared with the state-of-the-art interactive segmentation model, i.e., MedSAM, SAT achieves superior performance, scalability and robustness with text as prompts. SAT-Pro can also act as a foundation segmentation model, adapted to specific tasks for improved performance. We further demonstrate the usage of SAT as a powerful out-of-the-box agent for large language models, seamlessly providing grounding ability in clinical applications such as report generation. In summary, we believe our work can enhance versatile, efficient and easy-to-use medical segmentation in clinical applications, and contribute substantially to the development of generalist medcial artificial intelligence.
6 Code Availability
The code is available at https://github.com/zhaoziheng/SAT.
7 Data Availability of SAT-DS
The access to each dataset can be found in Tab. 4 and Tab. 5. The data process code to build SAT-DS and our train-test splits for reproducibility and benchmarking are available at https://github.com/zhaoziheng/SAT-DS.
Table 4: Download links of the 72 datasets in SAT-DS.
Dataset Download Link
AbdomenCT1K [57] https://github.com/JunMa11/AbdomenCT-1K
ACDC [9] https://humanheart-project.creatis.insa-lyon.fr/database/
AMOS CT [29] https://zenodo.org/records/7262581
AMOS MRI [29] https://zenodo.org/records/7262581
ATLASR2 [46] http://fcon_1000.projects.nitrc.org/indi/retro/atlas.html
ATLAS [72] https://atlas-challenge.u-bourgogne.fr
autoPET [16] https://wiki.cancerimagingarchive.net/pages/viewpage.action?pageId=93258287
Brain Atlas [81] http://brain-development.org/
BrainPTM [6] https://brainptm-2021.grand-challenge.org/
BraTS2023 GLI [60] https://www.synapse.org/#!Synapse:syn51514105
BraTS2023 MEN [36] https://www.synapse.org/#!Synapse:syn51514106
BraTS2023 MET [62] https://www.synapse.org/#!Synapse:syn51514107
BraTS2023 PED [32] https://www.synapse.org/#!Synapse:syn51514108
BraTS2023 SSA [3] https://www.synapse.org/#!Synapse:syn51514109
BTCV Abdomen [38] https://www.synapse.org/#!Synapse:syn3193805/wiki/217789
BTCV Cervix [39] https://www.synapse.org/#!Synapse:syn3193805/wiki/217790
CHAOS CT [31] https://chaos.grand-challenge.org/
CHAOS MRI [31] https://chaos.grand-challenge.org/
CMRxMotion [88] https://www.synapse.org/#!Synapse:syn28503327/files/
Couinaud [83] https://github.com/GLCUnet/dataset
COVID-19 CT Seg [55] https://github.com/JunMa11/COVID-19-CT-Seg-Benchmark
CrossMoDA2021 [14] https://crossmoda.grand-challenge.org/Data/
CT-ORG [75] https://wiki.cancerimagingarchive.net/pages/viewpage.action?pageId=61080890
CTPelvic1K [48] https://zenodo.org/record/4588403#ẎEyLq_0zaCo
DAP Atlas [28] https://github.com/alexanderjaus/AtlasDataset
FeTA2022 [67] https://feta.grand-challenge.org/data-download/
FLARE22 [56] https://flare22.grand-challenge.org/
FUMPE [59] https://www.kaggle.com/datasets/andrewmvd/pulmonary-embolism-in-ct-images
HAN Seg [69] https://zenodo.org/record/
HECKTOR2022 [4] https://hecktor.grand-challenge.org/Data/
INSTANCE [45] https://instance.grand-challenge.org/Dataset/
ISLES2022 [25] http://www.isles-challenge.org/
KiPA22 [22] https://kipa22.grand-challenge.org/dataset/
KiTS23 [1] https://github.com/neheller/kits23
LAScarQS2022 Task 1 [43] https://zmiclab.github.io/projects/lascarqs22/data.html
LAScarQS2022 Task 2 [43] https://zmiclab.github.io/projects/lascarqs22/data.html
LNDb [68] https://zenodo.org/record/7153205#Ẏz_oVHbMJPZ
Table 5: (Continued) Download links of the 72 datasets in SAT-DS.
Dataset Download Link
LUNA16 [82] https://luna16.grand-challenge.org/
MM-WHS CT [103] https://mega.nz/folder/UNMF2YYI#1cqJVzo4p_wESv9P_pc8uA
MM-WHS MR [103] https://mega.nz/folder/UNMF2YYI#1cqJVzo4p_wESv9P_pc8uA
MRSpineSeg [66] https://www.cg.informatik.uni-siegen.de/en/spine-segmentation-and-analysis
MSD Cardiac [5] http://medicaldecathlon.com/
MSD Colon [5] http://medicaldecathlon.com/
MSD HepaticVessel [5] http://medicaldecathlon.com/
MSD Hippocampus [5] http://medicaldecathlon.com/
MSD Liver [5] http://medicaldecathlon.com/
MSD Lung [5] http://medicaldecathlon.com/
MSD Pancreas [5] http://medicaldecathlon.com/
MSD Prostate [5] http://medicaldecathlon.com/
MSD Spleen [5] http://medicaldecathlon.com/
MyoPS2020 [70] https://mega.nz/folder/BRdnDISQ#FnCg9ykPlTWYe5hrRZxi-w
NSCLC [8] https://wiki.cancerimagingarchive.net/pages/viewpage.action?pageId=68551327
Pancreas CT [78] https://wiki.cancerimagingarchive.net/display/public/pancreas-ct
Parse2022 [50] https://parse2022.grand-challenge.org/Dataset/
PDDCA [74] https://www.imagenglab.com/newsite/pddca/
PROMISE12 [47] https://promise12.grand-challenge.org/Details/
SEGA [73] https://multicenteraorta.grand-challenge.org/data/
SegRap2023 Task1 [51] https://segrap2023.grand-challenge.org/
SegRap2023 Task2 [51] https://segrap2023.grand-challenge.org/
SegTHOR [37] https://competitions.codalab.org/competitions/21145#learn_the_details
SKI10 [41] https://ambellan.de/sharing/QjrntLwah
SLIVER07 [23] https://sliver07.grand-challenge.org/
ToothFairy [13] https://ditto.ing.unimore.it/toothfairy/
TotalSegmentator Cardiac [90] https://zenodo.org/record/6802614
TotalSegmentator Muscles [90] https://zenodo.org/record/6802614
TotalSegmentator Organs [90] https://zenodo.org/record/6802614
TotalSegmentator Ribs [90] https://zenodo.org/record/6802614
TotalSegmentator Vertebrae [90] https://zenodo.org/record/6802614
TotalSegmentator V2 [90] https://zenodo.org/record/6802614
VerSe [80] https://github.com/anjany/verse
WMH [35] https://wmh.isi.uu.nl/
WORD [52] https://github.com/HiLab-git/WORD